### v1

# ”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder

Abstract.
Obsessive-compulsive disorder (OCD) is a mental health condition significantly affecting people’s quality of life. Although OCD can be effectively treated with evidence-based therapy (e.g., exposure and response prevention), Managing OCD symptoms independently, as an indispensable part of successful treatment, remains challenging due to fear confrontation and lack of appropriate support. We aim to comprehensively understand the challenges and needs in OCD self-management from the perspectives of both people with OCD and OCD therapists. Through interviews with 10 participants with diverse OCD conditions and seven therapists, we characterized different OCD symptoms, typical triggering factors, strategies, technology use, and barriers both inside and outside of therapy. Our findings highlighted gaps between OCD self-management needs and currently available support. Building on these insights, we suggest in-situ self-management technologies for OCD, including personalized symptom tracking, in-situ interventions, and discuss how OCD-specific privacy and social needs can be fulfilled with technology and beyond.
mental health, OCD, Obsessive-Compulsive Disorder, Exposure and Response Prevention, Acceptance and Commitment Therapy, in-situ intervention
copyright: noneccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
1. Introduction
Obsessive-Compulsive Disorder (OCD) is a mental health disorder affecting approximately 2% of world population (Sasson et al., 1997). People with OCD experience uncontrollable, recurring thoughts (obsessions) usually triggered by specific events (triggers), and engage in repetitive behaviors (compulsions) to relief the anxiety caused by obsessions (of Mental Health, NIMH). Contrary to the common misconception that OCD is simply ’being clean’ or ’focused on order’ (Spencer and Carel, 2021), OCD symptoms are highly diverse (Hirschtritt et al., 2017) and can significantly affect quality of life especially when compulsions become severe enough to disrupt daily activities (Macy et al., 2013; Eisen et al., 2006; Subramaniam et al., 2013).
Coping with OCD symptoms independently in daily life is crucial but challenging as real-life triggers can be too overwhelming to confront during intense OCD episodes without suitable guidance (Lind et al., 2013; Hezel and Simpson, 2019; Wheaton and Chen, 2021; Ojalehto et al., 2020). Many individuals with OCD seek professional support through cognitive behavioral therapy (CBT), a type of psychotherapy that help people identify and change unhelpful thoughts and behavior patterns (Beck, 2011). Despite its effectiveness (Yule and Whittal, 2017; Blakey et al., 2017; Hirschtritt et al., 2017), over 50% of people with OCD lack access to the therapy treatment (García-Soriano et al., 2014; Ruscio et al., 2010; Chong et al., 2012; Mayerovitch et al., 2003). Moreover, many who complete the treatment continue to experience residual symptoms and struggle to apply the skills learned in therapy to their daily lives (Eddy et al., 2004). These challenges highlight the needs for effective self-management technology that supports people with OCD in-situ beyond traditional therapy.
Recent advances in AI and sensing techniques have shown great potential for in-situ self-management for OCD—through intelligent scene understanding and behavior recognition (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024; Liu et al., 2024), such technology can detect triggering events and compulsive behaviors, enabling timely, in-situ interventions to help people with OCD address obsessions and resist compulsions during OCD flare-ups (Burchard et al., 2022; Olbrich et al., 2016). Despite the potential, research on in-situ intervention technologies for OCD remains limited and primarily focuses on specific symptoms, including recognizing excessive hand-washing on a smartwatch (Burchard et al., 2022) and detecting excessive checking behaviors via a short moving distance over an extended period of time based on GPS location (Olbrich et al., 2016). Given the complex and heterogeneous nature of OCD (Hirschtritt et al., 2017), timely research is needed to deeply understand the diverse personal experiences of people with OCD, uncover their challenges and needs, and identify suitable intervention feedback, thus inspiring effective, generalizable technology design.
To fill this gap, we contribute the first comprehensive study with both people with OCD and OCD professionals. We first interviewed 10 participants with diverse OCD symptoms about their detailed experiences of OCD, symptom self-management strategies, persisting challenges, and desired technology to assist with self-management. To complement these personal insights, we further conducted an interview with seven mental health care providers who specialized in OCD to learn about their treatment strategies throughout different stages of OCD therapy and technology recommendations to facilitate symptom management outside of therapy context, guiding the feedback design for in-situ intervention technology for OCD self-management.
Our studies detailed OCD people’s individualized triggers and compulsions and identified typical properties that predominantly affect trigger intensity. We found consistency between self-management strategies used by people with OCD and those taught by OCD therapists, while also uncovering nuances in treatment strategies and use of technology. By combining insights from both studies, we highlight gaps in OCD self-management despite currently available resources. By triangulating the findings from both studies, we derive technology design implications from trigger and compulsion detection, in-situ feedback design, and privacy perspectives. Beyond in-situ self-management, some implications can also benefit OCD therapy, such as supporting out-of-session practice.
In summary, our research contributes: (1) the first study that thoroughly understand in-situ self-management strategies and needs from perspectives of both people with OCD and OCD professionals (2) rich qualitative data that reveals detailed personal experiences of people with OCD to inspire in-situ technology design (3) design implications that focused on input (OCD trigger and compulsion recognition) and output (feedback design) of in-situ intervention technology and OCD-specific privacy and social needs.
2. Related Work
2.1. Obsessive-Compulsive Disorder & Therapy Treatments
2.1.1. What is OCD
Obsessive-Compulsive Disorder occurs when a person is caught in a loop of obsessions and compulsions (American Psychiatric Association et al., 2013). Obsessions are repetitive, intrusive, and unwanted thoughts, urges, or mental images caused by certain triggers and induce distress feelings such as fear, anxiety, and disgust (Association, 2024). Compulsions are repetitive overt behaviors or covert mental rituals (e.g., repetitively reviewing events mentally to prevent harm) performed to relieve distress caused by obsessions (Olatunji et al., 2013); avoidance of people, places, or situations that trigger obsessions can also be a manifestation of compulsion (Association, 2024). People with OCD might feel temporary relief by conducting compulsive behaviors (e.g., excessive hand washing), but in no way doing these compulsions can bring them pleasure (of Mental Health, NIMH). In severe situations, obsessions and compulsions can take up hours of a person’s day and can affect family and social relationships (Koran, 2000). As a result, OCD poses barriers to both professional and personal life, leading to diminished quality of life (Bobes et al., 2001; Subramaniam et al., 2013; Macy et al., 2013).
OCD is highly heterogeneous. Yale-Brown Obsessive Compulsive Scale (Y-BOCS) Symptom Checklist (Goodman et al., 1989) organizes different OCD symptoms into eight obsession categories and seven compulsion categories (Goodman et al., 1989; McKay et al., 2004). Common obsession themes include fear of contamination, persistent doubting, violent or sexual intrusive thoughts. In response to these obsessions, the compulsive behaviors can be excessive hand washing, repeated checking (e.g., whether the oven is off), and repeated neutralizing thoughts (e.g., “I’m not a bad person”), respectively (Hirschtritt et al., 2017). Unlike obsessions and compulsions, the diversity of triggers has been overlooked in OCD research. Existing studies focus on generic triggers for specific symptoms (Mataix-Cols et al., 2009) or describe OCD episodes without clearly distinguishing triggers (Knapton, 2016). Since triggers are highly personalized and important in characterizing OCD symptoms (Baioui et al., 2013), it is crucial to further understand and categorize triggers by exploring OCD people’s personal experiences.
2.1.2. Common Therapy Treatments for OCD
To combat OCD, cognitive behavioral therapy (CBT) is the most effective method (Hirschtritt et al., 2017), focusing on identifying and changing patients’ unwanted thinking and behavioral patterns (Beck, 2011). CBT includes different forms, such as exposure and response prevention (ERP) therapy, cognitive therapy (CT), and acceptance and commitment therapy (ACT) (Björgvinsson et al., 2007).
ERP therapy has been widely used to facilitate behavioral changes (Blakey et al., 2017; Björgvinsson et al., 2007; Hirschtritt et al., 2017). It is a psychotherapy that gradually exposes individuals with OCD to triggers that provoke obsessions while simultaneously discouraging compulsions (Blakey et al., 2017; Hirschtritt et al., 2017). To decide the exposure tasks, an exposure hierarchy is built to cover the clients’ OCD triggers ranked based on the intensity of distress they cause. In between sessions, clients are also expected to complete exposure practices to achieve better outcome. Once the client completes the most difficult exposure successfully, the treatment can be concluded. However, ERP’s nature of ‘facing the fears’ can distress the clients and be time consuming (Patel et al., 2021), hence resulting in high dropout rate (Ong et al., 2016).
To facilitate cognitive change, cognitive therapy (CT) is conducted to correct unhelpful and dysfunctional beliefs that lead to obsessions. Strategies are commonly developed from the underlying characteristics of the dysfunctional beliefs, such as overestimation of threat. An example strategy can be showing the client that the catastrophic outcome is a consequence of many steps so its probability is much lower than they expect (Yule and Whittal, 2017). Prior research shows that CT can achieve comparable effectiveness to ERP (Yule and Whittal, 2017), and that CT is better tolerated than ERP with lower dropout rate (Yule and Whittal, 2017; Björgvinsson et al., 2007).
More recently, acceptance and commitment therapy (ACT), as a newly developed form of CBT, has been applied to the treatment of OCD and can significantly reduce OCD symptoms (Philip and Cherian, 2021; Evey and Steinman, 2023). Unlike traditional CBT (e.g., ERP), ACT encourages people to accept the presence of obsessions and the distress resulting from them, without attempting to change these feelings through compulsive behaviors (Smith et al., 2017). It aims to improve the ability of people with OCD to stay at the present, not attending to previous or future experiences (Smith et al., 2017; Philip and Cherian, 2021; Evey and Steinman, 2023). In ACT, people with OCD are guided to identify the values that are more important to them and encouraged to commit to behaviors that are consistent to these values. For example, a person with OCD who has identified spending time with family as a value may find excessive hand washing interfering this important event of life. Research has shown that combining ERP and ACT can further improve people’s ability to respond to more intense obsessions (Ong et al., 2020).
In summary, CBT has proven to be effective in addressing OCD symptoms from both obsession and compulsion aspects. However, it is not available 24/7 whenever people need support to manage their issues (Blakey et al., 2017). The homework between sessions (e.g., symptom self-monitoring, self-guided exposure practices)—as an important part of CBT—could be difficult for clients to comply, affecting the outcome of the therapy (Wheaton and Chen, 2021).
2.2. Technology for OCD Symptom Self-Management
To overcome the challenges of OCD symptom self-management, a variety of different forms of new technology have been developed and evaluated. Existing efforts have been focused on symptom and behavior monitoring(Khan et al., 2011; Frank et al., 2023), proactive skill building(Khan et al., 2011; Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010), and in-situ intervention (Burchard et al., 2022; Olbrich et al., 2016). We discuss prior work on self-symptom management following this taxonomy.
2.2.1. Symptom monitoring
Symptom monitoring is an effective approach to supporting self-management for various mental health conditions, such as anxiety disorder and depression (Kruzan et al., 2023), due to its potential of improving a person’s symptom (e.g., thoughts, behavior) awareness (Kruzan et al., 2023). In the context of OCD, prior work has explored two ways of self-monitoring, (1) self-report, and (2) passive sensing.
Ecological Momentary Assessment (EMA) — a research approach in clinical psychology to periodically prompting for people’s behaviors and experiences in their natural environments — has been broadly used to track and assess an OCD person’s symptoms (Frank et al., 2023; Brown et al., 2020; Tilley and Rees, 2014; Rupp et al., 2020), and demonstrated high acceptability, practicability, and representativeness (Frank et al., 2023). However, research has shown that overly relying on self-report approach can result in noisy (Frank et al., 2023) and missing data.
In contrast, passive sensing enables rich data collection without burdening users. Researchers have investigated passive symptom monitoring using various sensors (Thierfelder et al., 2022; Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022). For example, Thierfelder et al. (Thierfelder et al., 2022) monitored OCD-induced stress by measuring heart rate, hand activity and gaze fixation, and demonstrated that stress and anxiety detection is possible using multimodal sensors. Other studies measured sleep quality of people with OCD using actigraphy (Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022). While symptom monitoring can improve symptom awareness, prior work has primarily focused on monitoring users’ internal states instead of contextual information which is important in characterizing OCD symptoms (Llorens-Aguilar et al., 2021). In addition, without proper knowledge and coping skills for symptom management, as well as actionable insights derived from collected data, the effectiveness of symptom monitoring can be further limited.
2.2.2. Proactive Skill Building
To support out-of-session symptom management skill building, smartphone-based intervention has been developed that incorporate evidence-based therapy treatment in peoples’ everyday life (Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010). LiveOCDFree is an app that guides the user to set up and conduct self-administered ERP by creating an exposure hierarchy, and identifying specific exposures to practice (e.g. touching doorknobs). The user is able to set reminders for ERP practice and monitor their progress. An evaluation with 21 participants demonstrated significant improvement in OCD and anxiety symptoms after using the app (Boisseau et al., 2017). Similarly, NOCD supported self-administered ERP between therapy sessions incorporating ACT strategies. A study with 2,069 participants showed that NOCD effectively reduced OCD symptoms (Feusner et al., 2021). However, there are limitations with self-administered ERP such as inflexible template for specifying obsessions and compulsions which restricts user autonomy (Hong et al., 2020). Prior work has also reported clinical significance of web-based CBT in reducing OCD symptoms (Ferreri et al., 2019), such as BT Steps, a program that guides users with OCD to learn symptom management skills (Schure et al., 2020).
These technologies effectively supported OCD people’s symptom management outside of therapy. However, they all require OCD clients to proactively practice to build up coping skills, instead of providing targeted support for clients to respond to intense real life triggers, which can be difficult to manage without therapists’ help (Lind et al., 2013). In contrast, in-situ intervention technologies that implement therapeutic strategies (e.g., exposure practice) along with real life events (Kim et al., 2024) can potentially assist with OCD treatment in a more effective manner.
2.2.3. In-Situ Intervention
As opposed to pre-scheduled therapy sessions, the increasingly powerful mobile and sensing technologies allow for more flexible and intelligent mental well-being interventions. Such systems can recognize and adapt to users’ internal mental states and physical context in daily life to provide just-in-time mental support — an intervention design that provides the right type and amount of support, at the right time, in the context that the person needs it most and is most likely to be receptive(Nahum-Shani et al., 2018).
Such interventions have been applied to treat different mental health conditions, such as substance use disorder (Gustafson et al., 2014), depression (Burns et al., 2011), and schizophrenia (Ben-Zeev et al., 2014, 2013). For example, ACHESS (Gustafson et al., 2014) is a mobile application that tracks the GPS location of a user with alcohol use disorders and provides text-based support when they go to high-risk locations (e.g., a bar). An evaluation with 349 participants shows that ACHESS significantly reduced risky drinking days.
There has been limited exploration on in-situ intervention for OCD. Burchard et al. (Burchard et al., 2022) develops a deep-learning-based system to detect excessive hand-washing—a typical compulsion in contamination OCD—using the inertial motion sensor data on a smartwatch. However, although the system supports real-time symptom recognition, it does not provide in-situ feedback to the user. Olbrich et al. (Olbrich et al., 2016) develops a smartphone application Geo-Feedback that actively detects the user’s excessive checking behavior indicated by short moving distance over an extended period of time via GPS location and evokes a selected alarm when the excessive checking behavior is detected. A case study with a person with severe OCD indicates the effectiveness of the application. These examples have demonstrated initial success in designing real-time interventions for people with OCD. However, they are purely focused on physical compulsions and highly limited in context, hindering generalization to other OCD conditions. More research is needed to guide the design of in-situ intervention technology, including both the detection of OCD triggers and the design of effective and acceptable interventions consistent with therapeutic advice.
3. Study 1: Interviewing People with OCD
We first conducted semi-structured interviews with 10 participants with OCD. We aimed to comprehensively understand the impact of OCD (e.g., obsessions, compulsions, triggers) on their daily life, with an ultimate goal of distilling the needs and challenges of OCD symptom self-management and informing the design of in-situ intervention. The study was approved by Institutional Review Board (IRB) at our university.
3.1. Method
3.1.1. Participants
We recruited 10 participants (7 female and 3 male) whose ages ranged from 19 to 47 ($Mean$ = 28.5, $SD$ = 8.5). Participants’ demographic information and OCD diagnosis are detailed in Table 1. We recruited participants from online OCD support groups and the research email service at our university. Participants were screened via an email to (1) inform them of the interview content to ensure they would be comfortable participating in this study, and (2) check their eligibility. A participant was eligible if they were (1) at least 18 years old, and (2) diagnosed with mild to moderate OCD and were willing to disclose their OCD symptoms with us. We did not recruit people experiencing severe OCD at the time of the study, because our interview questions can be triggering without adequate ability to manage symptoms independently. Participants were compensated for \$20 per hour. At the time of the study, all participants except P6 (from India) were located in the United States.
ID Age/ Gender Severity Diagnosed Time OCD Types
P1 19/F Mild-to-moderate 1.5 years ago Contamination OCD, responsibility OCD
P2 21/F Mild 1.5 years ago Harm OCD, sexual OCD
P3 25/F Moderate 16 years ago Contamination OCD
P4 22/F Moderate 3 years ago Just-right OCD
P5 47/F Mild 25 years ago Responsibility OCD
P6 25/F Mild 6 years ago Contamination OCD, responsibility OCD
P7 25/F Moderate 1 year ago Contamination OCD, responsibility OCD
P8 33/M Moderate 6 years ago Contamination OCD, responsibility OCD
P9 33/M Moderate 11 years ago Responsibility OCD
P10 35/M Moderate 6 months ago Just-right OCD
Table 1. Demographic information and basic obsession types of the 10 participants living with OCD in Study 1.
Our participants demonstrated diverse OCD types (Table 1), covering most common obsessions and compulsions in the Y-BOCS Symptom Checklist (Goodman et al., 1989). Six (P1, P5, P6-9) reported to have responsibility OCD, whose obsessions were around not being responsible enough or fear of making mistakes. Common compulsions reported by participants included checking if the door is properly locked. Five participants (P1, P3, P6-P8) had contamination OCD (e.g., dirt, germ) with common compulsions being washing hands and cleaning. P4 and P10 had just-right OCD which manifested as being obsessed with symmetry or the position of objects. The corresponding compulsion would be fixing asymmetry (P4) and objects that are not where they should be (P10). P2 had harm OCD, having intrusive thoughts about harm occurring to herself and people around her with compulsion being rumination about the negative thoughts and avoidance of the trigger. She also experienced sex-related obsessive thoughts (sexual OCD) when interacting with men and avoided looking at them to ease her distress. Table A.1 detailed the OCD experiences reported by participants.
Our participant demographic was skewed towards women. However, it is consistent with the higher prevalence of OCD in adult females than adult males (Mathes et al., 2019). Similarly, consistent with the distribution of different OCD subtypes (Ball et al., 1996), more than half of our participants had contamination OCD and responsibility OCD.
3.1.2. Procedure.
The study was a single-session interview via Zoom that lasted 90 minutes. We first asked participants about their demographic information and diagnosis of OCD. Then we asked participants about their detailed personal experience with OCD. We prompted participants for their obsessions, triggers, and compulsions in different living environments (e.g., home, workplace, public spaces). For triggers, we specifically asked for finer-grained details with a goal of identifying the characteristics and rationales of triggers. Participants were prompted to describe strategies to manage their OCD symptoms either recommended by therapists or developed by themselves. We also asked about their therapy experiences and use of other resources (e.g., technologies) to assist with symptom management. Finally, participants discussed ideas about desired technological support for OCD symptom self-management.
3.1.3. Data recording and analysis.
We audio recorded and transcribed all interviews via an online transcription service (otter.ai [^1] https://otter.ai). One researcher manually cleaned all transcripts and corrected transcribing errors by checking the recorded videos. We analyzed the data using thematic analysis (Clarke et al., 2015; Braun and Clarke, 2006) We selected two representative transcripts as samples, and two researchers coded the samples independently with open coding. Then, they discussed and reconciled their codes to resolve any differences and developed an initial codebook upon agreement. One researcher then coded the remaining eight transcripts based on the codebook. If a new code emerged, the two researchers discussed on the code and updated the codebook upon agreement.
We derived themes using a hybrid of inductive and deductive approaches (Swain, 2018; Braun and Clarke, 2006). First, we generated high-level themes based on fundamental concepts of OCD symptoms and management (e.g., OCD symptom management strategies) (Wu and Lewin, 2017; Knapton, 2016) using a deductive approach (Crabtree and Miller, 1992). Next, we used an inductive approach (Boyatzis, 1998) to generate sub-themes and identify new themes (e.g., connecting triggers to compulsions) by grouping related codes via affinity diagramming (Clarke et al., 2015; Braun and Clarke, 2006). Two researchers then cross-referenced the original data, codebook, and themes to make final adjustment, ensuring all codes were grouped in the correct themes. Our analysis resulted in over 400 codes and 23 themes and sub-themes. We present our themes, subthemes and example codes in Table A.2.
3.1.4. Ethical considerations and reflexivity statement.
Our interviews involved discussing participants’ OCD symptoms, which could cause mental distress. To provide a safe and comfortable environment for participants, we consulted a clinical psychologist (a co-author) with experience treating OCD to ensure appropriate language was used in the interview. We applied people-first language (Sharif et al., 2022), such as people with OCD instead of OCD patients, to promote inclusion and reduce stigma. Participants were informed they could skip any uncomfortable questions and turn off their camera if preferred. We recognize that researchers’ identities and background may influence our data interpretation. The primary researcher identifies as a person with OCD. His background and lived experience with OCD enabled us to develop appropriate questioning routes, better empathize with participants, and interpret their responses via a unique lens. Moreover, the researcher team comprises diverse expertise, including HCI, accessibility, intelligent system design, and OCD therapy, enabling us to better contextualize the study data in existing technology and mental health literature and derive suitable technology implications to support people with OCD.
3.2. Findings
In this section, we first report participants’ detailed experience of OCD in their daily lives and coping strategies. We then unpack the challenges faced by participants despite currently available support.
3.2.1. What Makes Something a Trigger?
We first investigated the OCD triggers (Table A.1) to determine what to recognize and how to recognize triggers for self-management technology for OCD. With this goal in mind, we categorize OCD triggers into different types and examine the key attributes that contribute to a trigger, which has been overlooked in prior literature in OCD (See Section 2.1.1).
Spectrum of Triggers. OCD triggers are highly idiosyncratic (Baioui et al., 2013). While most triggers reported by participants are specific physical objects (i.e., triggering objects), such as trash cans for contamination OCD (P1), and knives for harm OCD (P2), participants revealed other types of triggers, including contexts, sounds, and even thoughts.
We found that OCD can be triggered by participants’ physical context, such as specific locations or times (P1, P3, P7-9). For instance, participants with contamination OCD (P1, P3, P8) were triggered every time they returned home, since germs brought from the outside can contaminate their home environment. Interestingly, weather can also be a trigger. As P7 commented, her fear of hitting a pedestrian emerged with low visibility on the road during rain or snow. OCD triggers can also be audible (P3, P6-8, P10). For example, coughing sound can immediately trigger P3’s contamination OCD, and accident-related keywords in a conversation would prompt P8 to check his own apartment for safety. In addition, some participants reported mental triggers (P5), such as unspecified worries, without clear links to physical objects. Moreover, OCD episodes could sometimes occur even without identifiable triggers (P1, P4, P6, P10).
Through a deeper investigation into triggers, we found that for object triggers, different properties of a trigger can significantly affect trigger intensity (i.e., the degree of distress caused by a specific trigger), rather than the object itself, leading to different levels of anxiety and compulsion. We summarize the following trigger properties across different types of obsessions.
Sensory Attributes. Visual attributes such as color can alter the intensity of a trigger (P3, P7). For example, P7’s contamination OCD were triggered when the dirt presents high contrast against its background, such as coffee grind on a white countertop. While the same amount of salt would not be as triggering. In addition to color, some specific shapes can also be more triggering than others, such as sharp edges. P2 felt more uncomfortable when interacting with objects that have sharp tips in her home environment, including knives, pencils, and the sharp corners of newel posts on a staircase. Moreover, tactile attributes of a triggering object is another factor that affects trigger intensity (P1, P7). The sticky texture of sauce (e.g., ketchup) caused more contamination concerns to P1 and P7 than solid food.
Affordances. We found triggering obejcts that afford direct contact with the user can cause higher level of distress to participants with contamination OCD (P1, P3, P6-8). For example, door handles in public spaces and handrails in transportation are common triggers because it is meant to be touched using hands by many people, thus considered dirtier than surfaces not requiring direct contact (e.g., automatic doors). P1 mentioned trash can as one of her OCD triggers, but the ones with a lid and handle were more triggering for the same aforementioned reason.
Spatial Attributes. Spatial attributes of a triggering object is another property that determines trigger intensity. First, the quantity of triggering obejcts can affect the intensity of contamination OCD. P3 tended to avoid occasions with big crowds of people more than others due to concerns about contagious diseases. Second, the proximity of a trigger to a person with OCD is another factor that alters trigger intensity. According to P2, a knife in hand caused more harm-related intrusive thoughts to her than a knife on the counter top. Interestingly, trigger intensity can depend on the pose and orientation of a physical trigger. P2’s sexual intrusive thoughts were more likely triggered when a man was in a position that “resembled those in her intrusive thoughts.”
Certainty of Contamination Origins. For participants with contamination OCD, trigger intensity was influenced by the perceived certainty of the contamination’s origin (P1, P8). For instance, P1 was not triggered by food residues on a plate if she knew who was responsible for the mess. Similarly, P8 was more concerned about germs brought by strangers to his apartment than by his own family, whom he trusted for their hygiene.
Other Factors Affecting Trigger Intensity. Six participants (P1-4, P6, P7) further pointed out that their OCD triggers intensified when experiencing mental stress. It can be from schoolwork (P1, P2, P4), unfamiliar environment (P1) and personal life (P4). Additionally, poor self-care such as sleep deprivation and physical inactivity is another factor that exacerbates trigger intensity (P1, P2). Interestingly, unlike mental stress, physical stress could sometimes help reduce trigger intensity (P7, P8). As P8 explained, “if I’m pretty tired, my brain can’t put so much energy towards the OCD.” Similarly, trigger intensity can be reduced when short-term memory is fully occupied by other tasks (P2, P6), e.g., when working (P6).
In summary, trigger intensity is affected by not only trigger itself, but also properties of triggers whose variation can result in significantly different anxiety levels. Since OCD symptoms are highly personalized (Hirschtritt et al., 2017), trigger properties should be accounted for when designing systems for trigger recognition. Moreover, given the diverse types of triggers beyond physical objects (e.g., sounds, contexts), multi-sensing techniques and multi-modal recognition methods should also be considered in trigger detection.
3.2.2. Connecting Triggers to Compulsions
Besides triggers, compulsion is another important indicator to trigger in-situ interventions for self-management technology. Unlike prior OCD literature that merely focused compulsions on user behaviors (e.g., washing hands) (Starcevic et al., 2011; Goodman et al., 1989), we found that our participants constantly associated their compulsion behaviors with triggers. We thus categorize compulsions into five categories based on their relationship with triggers to inform compulsion recognition for in-situ intervention technology, including: 1) direct trigger interaction, 2) consequence mitigation, 3) trigger avoidance, 4) assisted compulsions, and 5) mental compulsions. As trigger avoidance (e.g., avoiding contact with objects deemed dirty by people with contamination OCD) and mental compulsion (e.g., rumination) has been recognized in existing OCD research (Association, 2024; Williams et al., 2011), we focus on other three types of compulsions below.
Direct Trigger Interaction. Direct trigger interactions are physical behaviors that engage with the trigger to temporarily resolve the distress caused by the trigger. For participants with responsibility OCD, repeatedly examining the trigger is a common compulsion, such as checking or jiggling the door locks (P8, P9). Repeated checking was also observed in P3 who had contamination OCD. She would constantly check the color of the meat to ensure doneness while cooking. Ordering and organizing is another direct interaction with triggers to precisely control the position, orientation, or order of the triggers to avoid accidents (e.g., keep appliances unplugged when not using, P7, P8), reduce harmful thoughts (e.g., pointing the knife to herself to protect others, P2), and keep life organized and “just right” (e.g., having to park his car in the same spot, P10). Direct trigger interactions can also happen when the OCD trigger is a person’s body part. For example, P4 would pull her eyelash or hair so both sides look even whenever she examines her look in the mirror. However, while temporally relieving mental distress, such compulsion could bring long-term mental health issues, “[Hair and eyelash] takes the longest to grow back and it really impacts my confidence and stuff, because it’s noticeable.” (P4)
Consequence Mitigation. Consequence mitigation is a compulsive behavior that addresses the speculated aftermath after interacting with a physical trigger when it is inevitable. For participants with contamination OCD, changing/cleaning clothes when coming back home and washing body parts after contacting something ‘dirty’ are common compulsions. Additionally, P3 mentioned self-body scanning for potential symptoms when encountering someone appearing sick. For participants with responsibility OCD, repeated checking can also be consequence-oriented. For example, participants who worried about hitting pedestrians while driving would check for marks on the car later (P7) or look at rear mirror to confirm no person was run over (P9). P7 described this lengthy process, “It used to take me like 10 minutes to check all of my car because I would get out and I would check like the doors and the windows, and I would check the hood of the car and the back of the car for scratches.”
Assisted Compulsions. In addition to performing compulsions independently, some participants sought accommodations from others (P1, P3, P6-8, P10) to help them complete compulsions. For instance, P10 disclosed their OCD to gain understanding when their symptoms affected others. He described the struggle of resisting a compulsion and how he ultimately requested someone to move their car from his preferred parking spot, “I went and knocked on the person’s door and told them to move their car, as unreasonable as that sounds. At this moment, I physically couldn’t stop myself, I had to. It was a compulsion I could do nothing about… I’m sitting here sweating, crying… you know, almost in physical pain, over the fact that my car is not in its ‘correct’ spot. And then I just said, ‘Okay, I have to go talk to this person.’ ”
Some participants also proactively asked (P1, P8) or passively waited (P3, P7) for help from other people to share the burden of performing compulsions. For example, P8 asked his friends to reassure him that the faucet was off instead of checking it repeatedly, since he trusted them more than himself. P3 would feel relieved when someone held doors for her to avoid contamination.
Effect of Environment Familiarity on Compulsions. Interestingly, five participants argued that the length and frequency of their compulsions was affected by their familiarity with the environment. P6 had more compulsive behaviors when around closed people than strangers, and four other participants (P2, P4, P8, P10) experienced fewer compulsions in public space. P4 indicated that she did not want others to see her compulsions, while P6 and P10 ‘gave up’ on compulsions in public knowing that they have no control over the environment.
In summary, connecting triggers and context information to compulsions can potentially facilitate compulsion recognition by tracking not only user behaviors but also the relationship between users and the triggers or the environment.
3.2.3. OCD Symptom Management Strategies
To combat OCD, participants adopted various strategies from OCD therapy (P1-5, P7, P9, P10), support groups (P5), online education resources (P8), and personal experiences (P6, P10), to manage their OCD symptoms in different situations. We summarize participants’ coping strategies to inspire the feedback design of in-situ intervention for self-management technology.
Exposure Practice in Daily Life. To improve symptom management ability in a controlled environment, participants would do planned exposure practice as instructed by their therapists (P1-3, P5, P7, P10). Specifically, they exposed themselves to a self-created triggering environment at a scheduled time, and resisted compulsions according to the principles of ERP (details in Section 2.1.2), while monitoring their anxiety level. P3 articulated her exposure practice targeting the fear of touching batteries, “It (exposure practice) was holding a battery in my hand, rubbing the battery on my arms, putting it in my pocket, rubbing it on my clothes, my phone, on my bag… So I, essentially contaminating all these different things [with batteries] and then, not being able to wash my hands or do anything to decontaminate the battery acid… and having to sit with worrying that everything was contaminated.” Although the flexibility of adjusting trigger intensity allowed participants to gradually build up their tolerance to triggers, most triggers were unplanned in real life, and some triggers that induce intrusive thoughts could not be anticipated (P2).
Therefore, when confronted with unplanned triggers, participants would seize the opportunity to do a natural exposure practice. Due to the spontaneity of unplanned triggers and unpredictable trigger intensity, natural exposure practices were perceived to be more difficult (P1-P3, P7, P10). Despite the effectiveness of both planned and natural exposure practices, participants found them challenging especially at the beginning (P2) and tended to avoid exposures when alone (P1). They had to force themselves to complete the practices, since compulsions were too difficult to resist (P1-P3, P10). This reality surfaces a need for in-situ technological support that can assist people with OCD to address their obsessions and compulsions when encountering intense triggers.
Accepting & Challenging Obsessive Thoughts. Participants adopted different mental strategies to address their obsessions, adopted from acceptance and commitment therapy (ACT) (P2, P3, P5, P7-P9) and cognitive therapy (P1, P7, P8) (see Section 2.1.2). We summarize strategies used by participants below: (1) Accepting Uncertainty: participants used this strategy to accept the inherent uncertainty in life when their OCD was driven by a need for certainty. For example, P3 described her approach of coping with fear of batteries, “Maybe I will die because I am holding this battery. Maybe I won’t die. I don’t know. So a lot of kind of that. Maybe this will happen, maybe not. And sitting with that uncertainty.” (2) Refocusing on Values: participants identified their values (i.e., important aspects of life) and used them, rather than obsessions, to guide their choice of everyday behaviors. (3) Reinforcing Positive Thinking: participants re-evaluated their negative thoughts in a more accurate way to reduce obsessions. For example, P7 reframed her intrusive thoughts about unconsciously hitting pedestrians while driving by questioning the presence of evidence, such as any unusual sound.
Leveraging the self-reflective nature of these mental strategies, we can design conversation-based persuasive intervention (Orji and Moffatt, 2018) to guide people with OCD to adopt beneficial behaviors and thoughts when obsessions occur, such as intelligent agents that help users navigate their thought processes and address cognitive pitfalls.
Reducing the Urge for Compulsions. To suppress the urge of doing compulsions, both mental and physical strategies were used. Participants would re-evaluate the necessity of compulsions mentally (P2-4, P6). By reflecting on past OCD episodes, participants reminded themselves that nothing bad happened when compulsions were not done (P2) and that compulsions were never working (P4). Some weighed the harm compulsions caused against the relief they brought (P3, P4). For instance, frequent hand washing killed germs but caused dry, irritated skin (P3).
In addition to mental persuasion, participants also employed physical strategies to make compulsions more difficult to act on, e.g., P4 creatively put Vaseline on her eyelash and wore glasses to prevent herself touching her eyelash. When all these strategies failed to work, participants (P1, P3-P5, P8, P10) would use self-distraction to temporarily step away from their OCD by switching their attention to casual activities such as watching TV (P3), listening to music (P8), calling a friend (P1), and going out for a walk (P5).
In contrast to strategies aimed at suppressing the urge for compulsions in the moment, participants with responsibility OCD used reassurance logging to gradually reduce the length and frequency of compulsions (P8, P9). They commonly took pictures or videos of their checking behavior to simplify future checking behaviors. P8 articulated how reassurance logging alleviated his compulsions, “If I know I’m gonna be gone for a while. I will take a video where I’ll video my entire apartment, and then I’ll also turn the camera around and say, ‘I have checked the faucet this many times; all plugs are unplugged…’ so it’s almost like telling myself through a video in the moment that I’m good to go. And then if I ever get anxious, I can look back at the video.” Although some participants acknowledged that reassurance logging contradicts the principles of ERP (facing their fear), they believed it could complement ERP by gradually building confidence in one’s memory, which could ultimately reduce the need for compulsive checking (P9).
In summary, participants effectively reduced the urge for compulsions by drawing on past behaviors and memories of previous OCD episodes. Inspired by these strategies, symptom and behavior logging could be leveraged to design in-situ intervention methods that increase the cognitive burden of engaging in compulsions.
Relaxation. Unlike above strategies, relaxation techniques were used by participants to reduce the anxiety caused by intense triggers without seeking to break the obsession and compulsion cycle (P1, P3, P6, P7, P10). Such strategies included breathing techniques and muscle relaxation.
Self-Compassion. Besides in-situ strategies to combat obsessions, compulsions, and reduce distress, staying motivated is important in symptom self-management. When experiencing setback in symptom management that is usually caused by increased trigger intensity, participants would cheer themselves up (P2, P3) and acknowledge their achievements periodically (P5). P5 commonly noted down her reflection of success in symptom management—strategies and thought patterns that effectively addressed her obsessions or compulsions.
External Mental Health Support. In addition to self-management strategies, most participants received different forms of support from their family members and friends (P1-8, P10) to help manage their OCD symptoms. Some participants had family or friends who understood their mental health conditions and followed therapeutic advice (P2, P4, P7, P10). They helped participants monitor obsessions and compulsions to improve self-awareness (P4) and encouraged them to challenge their OCD independently (P1-3, P7). In contrast, some participants had family or friends who lacked an accurate understanding of OCD and did not want to see them “suffer.” As a result, they provided accommodation to ‘help’ them avoid triggers, which is usually against therapists’ suggestions. For example, P6’s mom washed her hands every time after taking elevators to accommodate P6’s contamination OCD.
Additionally, participants benefited from support groups (P3, P5, P6) and social media (P1, P8), where people with OCD share their experiences and symptom management strategies. To our surprises, none of our participants reported using any technologies dedicated to mental health support.
3.2.4. Challenges in OCD Symptom Management
Despite currently available resources and strategies, participants still experience challenges in OCD management.
OCD-Induced Social Barriers. Participants discussed the negative impact of OCD on their social life. Due to the heterogeneity of participants’ OCD symptoms, understanding the nuanced experience of OCD could be difficult, even with OCD professionals (P6, P10). P10 had to downplay his symptoms to gain empathy from his therapist.
Outside of therapy, participants often felt their compulsions interrupted social interactions (P2, P7, P8, P9). For instance, P9 found it embarrassing to go back to check the parking brake of his car during group activities. Moreover, exposure practices in real life can also bring social barriers since practicing in public spaces or when other people are around could be socially awkward (P1, P7).
Inconvenient Symptom Logging & Tracking. Daily symptom tracking and logging is crucial in helping participants improve self-awareness of their behaviors and thoughts (P1, P4, P5, P8). Participants typically recorded daily triggers (P4, P8), their responses to those triggers (P1, P4), and distress level during exposure practices (P1, P5) either physically or digitally (e.g., Google Doc). However, participants complained about the limitation of traditional journaling methods since they are unable to provide insights into participants’ symptom patterns. When asked about desired technology to close the gap, they proposed an intelligent tracking technology that could recognize their symptom patterns and recommend optimal management strategies based on logging history. However, P7 raised concerns over the privacy and emotional risks of such technology, “I would be most concerned about privacy issues, because it’s very vulnerable for people… to do all of that exposure [practices] in a place where their data is [at stake]. So if sessions are recorded… or their speech or their notes are recorded, making sure all of that is well protected… ”
3.3. Discussion
Combating OCD is not an easy task, even with support from various sources. While most participants benefited from psychotherapy and developed coping strategies based on their lived experience, tracking and managing OCD outside of therapy sessions can still be challenging. Many of these strategies were adapted from psychotherapy, however, without mental health professionals’ feedback, it is unknown whether they were effectively implemented in practices. To complement OCD participants’ experiences and guide the design of future intervention technologies, it is crucial to consider OCD professionals’ recommendations for effective symptom management, and further identify gaps between these recommendations and the practices of those living with OCD.
4. Study 2: Interviewing OCD Therapists
To gain insights on effective OCD symptom management from OCD experts, we conducted semi-structured interviews with seven mental health professionals who specialized in OCD. The study was approved by the Institutional Review Board (IRB) at our university.
4.1. Method
4.1.1. Participants.
We recruited seven OCD professionals (5 female, 2 male). Their age ranged from 29 to 60 ($Mean$ = 43.9, $SD$ = 10.4). Four participants (T1, T2, T4, T6) had more than 10 years of experiences working as mental health providers. Their specialization ranged from OCD, anxiety disorders, depression, eating disorders, PTSD, to substance use disorder. Participants had experiences in treating OCD across various severity levels (from mild to severe) and in different practice modalities (both hybrid and virtual). Note that the gender imbalance in the participant demographic is consistent with the reality that mental health therapy providers are predominantly women in the US (Clay, 2017; of Labor Statistics, 2023). Table 2 shows participants’ detailed information.
We emailed recruitment information to local OCD therapists listed on International OCD Foundation, a nonprofit organization providing resources and education for people with OCD (IOCDF, 2024). Participants were eligible if they were over 18 and had experience conducting OCD related therapy. We limited the recruitment to individuals who spoke English and located in the United States. Upon completion, each participant received \$40/hour as compensation.
ID Age/ Gender Certificate Years of Experience Specialization OCD Clients Severity Level Practice Modality
T1 46/F Licensed Clinical Social Worker 25 years OCD, anxiety disorders, eating disorders, PTSD, depression, substance use disorder Mild to severe Hybrid
T2 47/F Clinical Psychologist 14 years OCD, axiety disorder Mostly moderate to severe Virtual
T3 36/M License Professional Counselor 9 years OCD, anxiety disorder, phobias, health anxiety Mild to severe, mostly moderate Virtual
T4 60/F Licensed Clinical Social Worker 13 years OCD, anxiety, depression Mostly moderate Hybrid
T5 51/F Professional Counselor Training License (LPC-IT) 2 years and 4 months OCD, anxiety disorder, depression Mild to severe, mostly moderate Hybrid
T6 38/M License Professional Counselor 17 years OCD, anxiety disorders, eating disorders, PTSD Mild to severe Hybrid
T7 29/F License Professional Counselor 2 years OCD, anxiety disorder Moderate to severe Virtual
Table 2. Demographic information and mental health experience of the 7 OCD professionals in Study 2.
4.1.2. Procedure.
All interviews were conducted via Zoom. Each interview consists of three phases that overall lasted about 90 minutes. The first phase focused on participants’ demographic information, certificate, and experience as a mental health care provider. In the second phase, we asked about participants’ strategies and technologies used (if applicable) in different treatment stages. If a participant conducted ERP, we further asked for details about exposure creation and progress monitoring during exposure practices. We then asked about out-of-session homework and external support clients might need. In the last phase, we prompted for the challenges of current therapy practices as well as their opinions on current technological support for OCD.
4.1.3. Data Analysis.
Similar to Study 1, we conducted thematic analysis on interview transcripts. Two researchers coded two representative transcripts (samples) independently and generated a initial codebook. One researcher coded the remaining five transcripts using the codebook, updating it with new codes after discussion and agreement with the other researcher. Following a hybrid approach (Swain, 2018; Braun and Clarke, 2006), we first generated themes deductively based on therapy treatment process and challenges identified in prior work (e.g., psychoeducation) (Blakey et al., 2017; Wu and Lewin, 2017; Yule and Whittal, 2017). We then generate sub-themes and themes (e.g. using relaxation techniques suitably) inductively (Boyatzis, 1998). The final codebook contained over 300 codes, which we grouped into 16 themes and sub-themes (Table A.3).
4.2. Findings
4.2.1. OCD Diagnosis and Assessment
In the initial assessment of a client’s OCD symptoms, all participants used standardized self symptom rating tools, such as Y-BOCS (Goodman et al., 1989), to gain a basic understanding of their obsessions and compulsions. Since compulsions can often manifest in ways clients might not even realize, such as trigger avoidance and mental compulsion (e.g., rumination), our therapists usually conducted interviews to identify the presence of such mental compulsions (T1-T3, T5, T7). Due to clients’ lack of awareness of avoidance or mental compulsions, identifying them in early stage is difficult but important for treatment.
Participants developed criteria to identify compulsive behaviors and assess their impact on daily life:
  • Functional Impairment (T1): Evaluating how a behavior limits essential activities, such as avoiding social interactions due to intense mental compulsions (e.g., avoiding going outside).
  • Behavioral Rationale (T5-T7): Determining whether the behavior serves a specific purpose, or if the goal is merely to reduce anxiety (e.g., frequently seeking reassurance from others).
  • Common Sense (T7): Using common sense to assess whether a behavior is compulsion by considering what others would do to achieve the same goal (e.g., excessive hand washing).
In addition to in-session assessment, homework was assigned for clients to self-monitor their OCD symptoms now that they have better awareness of their OCD (T1-T3, T5-T7). Overall, these strategies could potentially be integrated into self-management technologies to help improve OCD people’s symptoms self-awareness.
4.2.2. Psychoeducation to Gain Insights
A person’s insight in OCD refers to the degree to which one recognizes the irrational and excessive nature of their OCD symptoms (Wu and Lewin, 2017), which is considered to be a key factors for successful treatment (T1, T4, T6, T7). Clients usually gain insights through psychoeducation along with the initial assessment, where they were taught about the definition of OCD, how OCD therapy works, the expected outcome of the therapy (T1-T7). Psychoeducation happens not only didactically but also throughout the treatment process (T1-T4, T6, T7). Clients gradually gain insights by discussing their symptoms during therapy sessions and symptom self-tracking between sessions.
Moreover, to assist with in-session treatment, most participants believed that psychoeducation for clients’ family is equally necessary (T1-5) because clients’ family need to learn how to respond to their loved ones’ behaviors in a therapeutic and helpful way instead of accommodating their OCD (e.g., providing reassurance) (T2-4).
As such, psychoeducation should also be considered to incorporate into self-management technology for both OCD clients and their families to enhance insights and awareness.
4.2.3. Active Treatment via ERP
All participants conducted ERP for their clients. In this section, we reported participants’ strategies on identifying trigger intensity, creating exposures, and guiding clients through exposure practices while monitoring their symptoms and progress.
Building Exposure Hierarchy Using Subjective Experience. Participants usually built an exposure hierarchy with clients—a list of OCD triggers ranked based on the intensity of distress they induce—to prepare for the subsequent ERP therapy treatment. Along with initial assessment, participants identified triggers via interviews, self symptom rating tools, and the self symptom monitoring homework (T1-T5, T7) and rated the intensity of each trigger. Most participants adopted the Subjective Units of Distress Scale (SUDS) (Miegel et al., 2022), a 100-point Likert scale (T1), for self distress assessment. Clients were asked to assign each trigger a SUDS score to finalize the exposure hierarchy. Uniquely, T3 asked clients to rate their urge to perform compulsions instead of using SUDS to distract them from focusing on the distress associated with triggers, “ If just tracking SUDS of how anxious you feel now, I mean, that’s what people are doing already. They’re already hyper vigilant about their anxiety. So I need them thinking about something else other than how anxious they feel.”
Furthermore, similar to our findings in Study 1, some participants examined the properties of triggers that can make a certain trigger less or more intense to build a finer-grained hierarchy (T2, T3, T5), as T2 elucidated, “ There is continuing assessment that’s happening during that hierarchy building. Because then I’m asking questions like, what would make [this trigger] easier. What would make it harder? Maybe they’re just telling me like, ‘well… it depends on if someone’s with me, or if I’m by myself; it depends on how many floors there are in the building…’ whatever variables are important to the exposure.”
Designing Diverse Exposures. Based on the exposure hierarchy, participants designed various types of exposures, including in vivo exposures (T1-T5, T7), imaginal exposures (T2-T7), and material exposures (T2, T3, T5, T6). In an in vivo exposure, the client is expected to confront the trigger in real world while supervised by the therapist. For exposures that are difficult to create in real world, imaginal exposures were used where participants created a script together with the client and guide them to imagine a scenario. Although imaginal exposure is less triggering, it can be effective when used appropriately. Two participants (T2, T3) used ChatGPT (OpenAI, 2024) to generate scripts for imaginal exposure. For instance, T2 wrote an obituary using ChatGPT for a client who worried about death. In addition, material exposures, such as videos, images of triggers, and VR exposure, were typically used to bridge the gap between an imaginal exposure and in vivo exposure.
Exposure Practices. Participants guided clients to participate in exposure practices based on the exposure hierarchy, starting from low intensity triggers. Before clients are exposed to triggers, participants first teach them strategies to tolerate anxiety without doing compulsions, aligning with our findings in Study 1. During the exposure task, participants would continue to monitor clients’ SUDS level or urge for compulsions (T3) by periodic polling. Typically participants ended exposure practices when clients’ SUDS decreases (T2, T4, T7).
Besides using SUDS, participants detected clients’ distress level by observing their physical behaviors (T3, T7), such as fidgeting, head shaking, tight chest, staring off the trigger, or sometimes an outward expression of distress. Whenever these happened, participants brought them up in the discussion after the exposure practice to learn what happened. In addition to inhibiting cognitive compulsions, participants also identified clients’ engagement of mental compulsion during the exposure by rating their mental effort (T2).
During high intensity exposure practices where clients could be too afraid to continue, participants would dynamically adjust the exposure intensity (T1-T7). Some participants allowed milder compulsions compared to clients’ usual routines, such as compulsions with shorter duration (e.g., washing hands for shorter time, T3, T7) and lower frequency (e.g., less frequent checking, T2). Additionally, participants decomposed the exposure task into simpler steps (T1, T4). For example, T1 demonstrated an example where a client avoided opening a new book due to fear of contaminating the book with fingerprints. Thus instead of asking them to open the book right away, the client was asked to start with a simpler exposure task—taking the book off the shelf.
Issues with ERP. Overall, all participants were confident about the effectiveness of ERP. However, ERP was less effective in managing mental compulsions (T1, T2). T1 emphasized the limitation of ERP when clients compensated for their behavioral change with mental compulsions, “ They (clients) are doing mental compulsions to sort of cancel out the fact that they didn’t do the physical compulsion, like, okay, I didn’t [check] that doorknob. Okay, I’m making my therapist happy, but what I’m doing in my mind is, I’m going over my mantra 15 times, because that makes me safe. ”
Our findings revealed the thorough process of in-session ERP, from building exposure hierarchy, to exposure design and ad-hoc intensity adjustment, providing insights into designing personalized and flexible in-situ technological support for exposure practices.
4.2.4. Making Long-Term Progress
To maintain long-term progress, professionals highlighted suitable ways to use certain strategies, assigned homework outside therapy sessions, and recommended self-management technologies.
Using Relaxation Techniques Suitably. Participants agreed that relaxation techniques should be taught to clients with high anxiety level, but advised using them only after exposure practices, as clients are expected to sit with their anxiety rather than resolve it immediately (T1, T2, T5, T6). If not used correctly, relaxation techniques can become a new compulsion (T1, T2, T5). As a result, T7 chose not to teach relaxation techniques because they might exaggerate clients’ anxiety if they fail to work, “At one point, [the client] may be really distressed. And they’ve been practicing these breathing techniques that are not going to work for you, then it’s even more distressing, because the only thing you know how to do [with your OCD] is not gonna work.”
Promoting Out-of-Session Success via Homework. Homework is crucial for successful OCD treatment. Participants assigned homework for clients to practice both planned and natural exposures (T1-7) outside of sessions (Section 3.2.3), where clients were asked to record their compulsion duration, anxiety level, and time needed for anxiety to decrease during each practice (T2-T7). Along with exposure practices, clients were instructed to implement learned strategies to promote cognitive change, such as negative thought reframing (T5, T7), ACT-related strategies (e.g., uncertainty mindset) (T1-3, T5-7), which overlap with self-management strategies mentioned in Study 1.
By tracking their compulsive behavior and anxiety level, clients became more aware of their progress and achievements across therapy sessions (e.g., spending less time washing hands). Such positive reflections on progress can enhance confidence and perseverance (T2-4, T6). However, participants also recognized that self-tracking can be difficult to maintain, due to forgetfulness and other daily responsibilities (T3, T6). In addition, exposure practice itself can be overwhelming with multiple triggers present out-of-session, leading to low homework compliance (T3, T7). To address these issues, participants would identify specific barriers faced by individual client, such as fear of exposure (T1, T3), and adjust the homework content accordingly. Although monitoring clients’ progress seemed impossible out-of-session, participants sometimes had clients fill out exposure tracking forms and set up phone reminders to ensure homework completion (T3).
Technology Recommendation for Self-Symptom Management. To ensure clients implementing therapeutic advice outside of therapy, participants also recommended tools and technologies to help clients manage OCD at home (T2-5, T7). Dedicated OCD symptom tracking apps, such as OCDfeat (OCDfeat, 2024), were suggested to easily record compulsions and exposure homework (T2). They also recommended meditation apps, such as Calm (Calm, 2024), for clients to practice mindfulness (T5) and listen to the bilateral stimulation playlist to calm the nervous system (T4) (Kohl et al., 2014). These tools are intended for consistent out-of-session use to monitor symptoms and reinforce skills learned from the therapy.
Our participants also recommended apps for in-situ use in response to OCD episodes upon triggered. To reduce reassurance seeking behavior using specific apps (e.g., web browser), T3 suggested app redirection tool (e.g., ScreenZen (ScreenZen, 2024)) to limit access to those apps on clients’ phones. Interestingly, he also recommended text replacement feature to clients to prevent them from searching specific keywords for obsession-related reassurance, “ I’ll go in and have people create shortcuts where… if they’re going to Google ‘heart attack,’ it actually will try to replace ‘heart [attack]’ with like ‘puppies’ or ‘cute cat videos,’ or something like that.”
However, participants found low engagement in recommended technologies among clients, especially for tracking purpose (T2, T3). While T3 attributed the issue to poor usability of the outdated interface, T2 identified privacy concerns for clients whose OCD symptoms involve taboo thoughts. As she explained “ They (clients) do worry about putting that (taboo thoughts), even writing it down on paper… so they need a lot of reassurance about where that data is being stored and who has access to it.”
Although most participants prefer learning and recommending new technology to their clients, some were reluctant to introduce new tools in the fear of developing reliance on them (T1, T5). T1 expressed her concerns, “You want to encourage people to instead learn how to cope with things by themselves… right? That’s what we ultimately want. So maybe you could have [the technology] where it was only like the first month or the first 3 months, or whatever, and then you have a gradual step down, that might work… because [relying on technology] is counter-therapeutic. We want people to be able to help themselves.”
When talking about self-administered ERP via mobile apps (e.g., NOCD (NOCD, 2024)), participants recognized its merit in supporting users with low severity (T1-4). However, participants emphasized the importance of psychoeducation in OCD and acquiring adequate symptom management ability via in-person sessions before using such self-help tools (T1, T3, T4, T6, T7).
In general, participants highlighted the need for a tracking tool that preserves privacy. Inspired by participants’ strategies to facilitate long-term success, future self-management technology should guide users to choose strategies wisely, and incorporate progress reflection to maintain high motivation.
5. Discussion
As an important part of treatment process, self-management of OCD presents significant challenges. The needs for technological support has long been under explored. Through our two studies with people with OCD and the OCD professionals, we sought to triangulate the current gaps and opportunities in OCD symptom self-management. From both studies, we found that people with OCD employed strategies learned from mainstream evidence-based therapies such as ERP and ACT and developed experience-based strategies to resist the urge of compulsions, aligning with professionals’ advice. This confirms the efficacy and generalizability of strategies in OCD therapy.
However, the two studies also highlighted gaps in OCD self-management outside of therapy despite currently-available resources. First, while natural exposure practice is encouraged due to their abundant presence in people’s daily life, these practices were often difficult to complete and tended to be avoided, confirming prior research (Seibell and Hollander, 2014; Eddy et al., 2004). Besides the difficulty of facing the fear, identifying mental compulsions could be challenging and required adequate insights. The absence of physical compulsion does not guarantee the absence of mental compulsions despite going through ERP. Second, some experience-based strategies adopted by people with OCD (Study 1) were identified to be counter-therapeutic and suggested to be prohibited (e.g. reassurance) or used cautiously (e.g., relaxation techniques) by the professionals due to the potential of developing reliance (Study 2). Third, both OCD clients and therapists recognized the importance of symptom and activity tracking, however, traditional tracking tools present limitations and raise privacy concerns due to vulnerability and shame associated with OCD, resulting in the low technology adoption in Study 1 and echoing OCD therapists’ observations in Study 2. Finally, people with OCD faced persisting challenges in OCD self-management despite receiving therapy, highlighting the need for additional support to facilitate smooth social interactions.
The insights from both studies emphasized the needs for self-management technologies beyond formal therapy sessions. By triangulating the insights from people with OCD and the professionals, we identify diverse OCD triggers and properties, compulsive behaviors in relation to triggers, and various effective strategies to inspire the input (e.g., trigger and compulsion recognition) and output (i.e., feedback design) for in-situ self-management technologies for OCD. Furthermore, given the persisting challenges experienced by OCD people, we discuss how technology can address OCD-specific privacy issues and opportunities to support OCD people’s social needs.
5.1. Enhancing Symptom Recognition & Tracking with User Input
As an important step of insight establishing and symptom self-management, our studies underscored a need for supporting personalized, automatic, ad fine-grained tracking of OCD symptoms and exposure practices.
5.1.1. Supporting Personalized & Automatic OCD Tracking
From Study 1, we found that OCD triggers are highly individualized, taking diverse forms (e.g., objects, contexts, sounds, thoughts), and that their properties significantly influence trigger intensity. While symptom tracking is essential for improving insights, it is difficult to maintain without therapist supervision (Study 2). Therefore, future technology should incorporate OCD users’ input, allowing specification of triggers to support personalized and automatic trigger detection. Similarly, the system can prompt for the user’s compulsive behaviors based on different types of compulsions (Section 3.2.2) and criteria proposed by OCD therapists (Section 4.2.1). Recent advances in AI have made it possible to adapt pre-trained models to specific domains through few-shot learning (Parnami and Lee, 2022; Zhang et al., 2023). By leveraging just a few examples provided by users, these AI models can be personalized to effectively detect and track OCD symptoms within each user’s unique context. Given the varied form of triggers and compulsions (e.g., direct trigger interaction, consequence mitigation), multi-modal AI models (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024) can be used together with wearable devices (e.g., egocentric camera and microphone (Liu et al., 2024)) to identify and track triggers and compulsions beyond the scope of visual-based methods. Despite the merit of AI-based symptom tracking automation, false positive prediction results can potentially exacerbate OCD symptoms, since people with OCD often have low confidence in their memories (Radomsky et al., 2014). To minimize this risk, prediction models should communicate potential errors and involve users in assessing the decision-making process using explainable AI techniques (Hossain et al., 2023).
5.1.2. Fine-Grained Symptom Formulation
Besides highly personalized triggers, our findings suggest the importance of considering different properties of triggers to enable fine-grained exposure hierarchy building and symptom tracking. In complement to personalized trigger specification, future tracking technology should prompt the user to reflect on how trigger intensity is affected by different properties of that trigger (Section 3.2.1), measured with SUDS or urge to perform compulsion (Section 4.2.3). As such, an exposure hierarchy with fine-grained and structured trigger properties can be built. During tracking, the system can log fine-grained details of triggers and situate them in the hierarchy as well as identify physical compulsive behaviors, to better support the tracking of exposure practices.
Figure 1a shows a usage scenario that can be achieved with proposed system: the user identified knife as a trigger of his harm OCD, they then are prompted to answer how each property can affect the trigger intensity, for example, a knife can be more triggering when the knife is closer to him, so they assign ‘knife in hand’ with the highest intensity level, and ‘knife on countertop’ with a moderate intensity level. They are also prompted to describe their compulsion—‘avoiding the knife.’ The system will then detect the physical triggers and compulsions in real-time from the camera feed.
For mental compulsions which are difficult to recognize, potential technology can incorporate physiological sensing (e.g., EEG, ECG, eye movement) to detect mental stress (Xia et al., 2019; Al-Jumaily et al., 2021) and mental workload (Fan et al., 2020; Wang et al., 2023), inspired by OCD therapists’ strategies in Study 2. Prior research has shown great potential of using physiological sensing as a therapeutic approach to detect users’ internal states for various mental health conditions. For example, estimating the anxiety level during exposure therapy for anxiety disorder (Gazi et al., 2021; Šalkevicius et al., 2019), and detecting the presence of depression to aid clinical diagnosis(Saeedi et al., 2021; Ay et al., 2019; Thoduparambil et al., 2020). These methods can be adopted to support OCD self-management. However, context information should be incorporated with physiological data since both our work and prior literature (Llorens-Aguilar et al., 2021) indicated the importance of context in recognizing OCD symptoms.
Overall, the proposed technology can potentially improve tracking compliance and precision by formulating fine-grained exposure hierarchy through natural language, and automating OCD symptom tracking using state-of-the-art AI systems. However, we do recognize that our proposed technology could not capture the same level of detail for tracking mental triggers (e.g., worries).
5.2. In-Situ Feedback Design for OCD Self-Management
Consistent with prior literature (Lind et al., 2013; Hezel and Simpson, 2019; Ojalehto et al., 2020), exposure practices can be difficult to follow through outside of therapy sessions, due to the existence of fearful triggers in diverse context. Triggers can appear anywhere anytime and provoke users’ OCD spontaneously, the burden of dealing with unpredicted triggers and resisting compulsions thus all fall onto the users with OCD, potentially resulting in severe consequences. We thus propose the following technologies that provide in-situ intervention during high intensity exposures in people’s everyday life.
5.2.1. In-Situ Compulsion Prevention
Inspired by participants’ strategies in Study 1, and OCD therapists’ technology recommendation, future intervention technology should consider using restrictive interactions to make compulsions more difficult to perform. In the field of human-computer interaction (HCI), researchers have explored the design of such techniques to prevent smartphone overuse (Orzikulova et al., 2024; Lu et al., 2024; Xu et al., 2022), such as delaying a user’s tap for a certain period before it takes effect (Lu et al., 2024). In OCD context, these restrictive interactions can be adapted to help reducing compulsions in digital space, e.g., seeking reassurance online. For compulsive behaviors that happen in physical world, potential technology can explore the design of mixed reality (MR) -based interaction to suppress the urge for compulsions. Researchers has explored the usage of virtual barriers to ensure user safety in MR (Hoang et al., 2022; Sousa et al., 2019). Such design can be adapted for compulsion prevention. Since most overt compulsions involve visual examining, disturbing the user’s vision can effectively interrupt compulsions.
Figure 1. (a) Fine-Grained Symptom Formulation: an example of the user specifying triggers and trigger properties to build fine-grained exposure hierarchy; (b) AR Compulsion Intervention: an example of using visual augmentation (blurry effect) to disturb the user’s vision.
Figure 1b describes a scenario where the user with contamination OCD washes their hands after contacting triggers (e.g., dirty dishes). Once the system detects excessive hand washing behaviors (longer than 20s), it will blur out or render distracting patterns on the users’ hands and faucet — objects involved in hand washing compulsions. As such, the user’s vision becomes too low to examine their hands, and eventually stopped washing them. Adopting the idea of allowing milder compulsion (Section 4.2.3) when exposure becomes too intense, the invasiveness (e.g., blurry effect intensity) of the visual augmentation should be adjustable. Such intervention can be promising for direct trigger interaction and consequence mitigation that involves active interaction between the user and surrounding physical objects.
For trigger avoidance type of compulsions, such as avoiding eye contact with people that induce intrusive thoughts, the system can render visual augmentations that encourages interactions with the trigger. For instance, when avoidance of triggers is detected, the system can render a gray-scale filter on everything the user sees except the triggering person to encourage eye contact, and turn off the filter when eye contact is maintained for a preset length of time as an exposure practice.
Despite the benefits of restrictive interactions in compulsion prevention, overly restrictive interactions might lead to low user compliance and even backfire (Lu et al., 2024). Future work should explore the optimal balance between effectiveness and restrictiveness of such technologies in OCD context.
5.2.2. In-Situ Cognitive Support
Besides physical methods, mental support inspired by ACT and cognitive therapy is another important strategy that effectively adjust people’s attitudes on their obsessions. Self-managment technologies should thus consider employing contextualized suggestions to facilitate acceptance of thoughts, once mental distress or compulsion is detected. For instance, the technology can guide the user to acknowledge the uncertainty to the user’s feared consequences after encountering the trigger. Combined with professional expertise, recent advancement in LLMs have enabled such context-aware intervention to support general mental well-being (Kumar et al., 2024; Nie et al., 2024). Moreover, value-based intervention has been suggested and implemented in prior work (Khan et al., 2011; Xu et al., 2022) as a ‘wake up call’ to remind the user of committing to their values. For example, when the user with contamination OCD is compulsively washing hands, a photo of their family will be displayed on their phone to encourage them to stop, if they identified ’spending time with family’ as a value. Alternatively, we can design in-situ cognitive support based on cognitive therapy. Such technology can guide the user via conversations to challenge their obsessions, such as reframing obsessive thoughts to more hopeful ones (Sharma et al., 2023), and analyzing the likelihood of feared consequences to reject obsessive thoughts (Yule and Whittal, 2017). Although the feedback design was proposed for in-situ self-management, we believe they can be effectively adopted for out-of-session practices to reinforce therapy outcome.
5.3. Addressing Privacy Needs for People with OCD
Privacy concerns have been discussed in prior literature about personal information sensing and tracking (Harari, 2020; Torous et al., 2018). To preserve users’ privacy, the tracking technology should explicitly inform them about what kind of data are being collected, what kind of data are inferred from collected data, and who has access to collected data (Harari, 2020). When it comes to AI models, on-device inference is preferred to minimize the risk of privacy invasion (Lee et al., 2019). However, due to technology limitation, many AI services require sending user data to remote servers for inference (OpenAI, 2024). In the context of OCD, researchers should pay attention to sensitive information that can exacerbates OCD users’ mental health issues. Users should be prompted to identify the type of sensitive data, such as audio or video involving fear, and taboo thoughts (e.g., sexual and violent thoughts). The system should then actively detect sensitive information in collected data and obfuscate them, such as blurring out sensitive elements in videos and photos (Zhang et al., 2024; Tseng et al., 2024), and reframing sensitive text to less disturbing alternatives (OpenAI, 2024) before sending them out for inference. The collection of biometric data (e.g., physiological data for mental state inference) can also pose privacy concerns due to potential unauthorized use of those data to infer personal identity (Marciano, 2019). Future OCD technology should adopt privacy-enhancing approaches such as feature transformation (Melzi et al., 2024; Nandakumar and Jain, 2015) to lower the risk.
5.4. Reducing Frictions in Social Interactions
Our studies suggest that OCD poses challenge to smooth social interactions due to the need for compulsions. Prior work also revealed that people with OCD experienced more interpersonal problems primarily due to certain worries and rumination than people without OCD (Solem et al., 2015). Due to the low public awareness of OCD, people with OCD usually need to endure the awkwardness caused by their symptoms during social interactions. Disclosing one’s OCD status can earn accommodation or understanding from people (P10 in Study 1). However, not all people with OCD are willing to disclose their mental health status, particularly when their OCD involves taboo thoughts (e.g., P2’s sexual OCD in Study 1). For those who are still building symptom management skills, additional support is needed to reduce frictions caused by OCD in everyday social interactions.
With current technology, researchers can explore the use of motivational interviewing via conversational agent (Miller and Rollnick, 2012; Oewel et al., 2024) to encourage the user to keep engaged with social interactions. Technology that delivers unobtrusive intervention for emotion regulation can also be used to address the mental distress in such situation, e.g., reducing anxiety level via haptic feedback simulating a low heart rate (Costa et al., 2016, 2019; Kelling et al., 2016). That said, technology can provide limited assistance to satisfy OCD people’s social needs without improved public awareness of OCD, and appropriate social support. More stakeholders in the field, including people with OCD, mental health professionals, policy makers, should be involved to enable a more inclusive and accessible environment for people with OCD.
5.5. Technology Adoption Concerns
The low adoption rate of technology by people with OCD revealed in both studies raises concerns about the practicality of our proposed technology design opportunities. Prior research in accessibility has pointed out key factors that affect the adoption of assistive technologies, including ease of use (Scherer, 2017; Phillips and Zhao, 1993), trust (Scherer, 2017), self-consciousness (Parette and Scherer, 2004; Scherer, 2017), and social acceptance (Shinohara and Wobbrock, 2011; Scherer, 2017). These factors all can potentially affect the adoption rate of suggested technology for OCD in-situ intervention. In addition, mixed-reality is still in its early stages (Cummings et al., 2023), and options for consumer level physiological sensing technology is limited. Given this situation, mainstream technology such as smartphone- and smartwatch-based interventions might remain the most practical solution due to their broad user base, high social acceptance (Shinohara and Wobbrock, 2011) and increasingly powerful sensors and processors that can support our proposed intervention designs (Dai et al., 2021; Incel and Bursa, 2023).
6. Limitations
Our research has several limitations. First, while participants in Study 1 represented diverse OCD themes, it is unclear whether our findings apply to less common OCD themes, e.g., relationship OCD. Future work should include a broader range of OCD themes to explore differences in management strategies and technology needs. Second, we recruited participants with mild to moderate OCD at the time of the study (Study 1). Although nine had experienced severe OCD before therapy, their self-management strategies and needs may vary with symptom severity. Hence, our findings should be generalized to people with severe OCD with caution.
7. Conclusion
Our research contributed the first in-depth investigation to understand OCD people’s specific experiences and their technology needs for symptom self-management. Through an interview study with 10 participants with diverse OCD symptoms, we identified their individualized triggers, compulsions, properties that affect trigger intensity, and revealed their self-management strategies. With another interview study with seven OCD therapists, we revealed the nuances in treatment strategies in complement to those used by people with OCD. Combining the two studies, we identified gaps in OCD self-management despite currently available resources, and discuss design opportunities for personalized symptom tracking, in-situ OCD intervention, and addressing privacy and social needs that are unique to OCD.
Appendix A Appendix
ID Obsessions Representative Triggers Corresponding Compulsions
P1 Coming home from outside Washing hands and clothes
Concern with contamination Seeing dirty dishes with food residual Cleaning the kitchen
Using trash cans Avoiding trash cans
Concern with not being told the truth No specific triggers Seeking reassurance whether someone is telling the truth
P2 Holding a knife in hand Pointing knife towards herself to protect others being harmed
Intrusive thoughts about harm occuring to herself and others Being around family or friends Rumination about why she had the intrusive thoughts
Avoiding eye contact with people who triggered her OCD
Intrusive thoughts about engaging in sexual activity with men in an uncomfortable way Being around other men Looking away from people who triggered her OCD
P3 Touching batteries (considered dirty) Washing hands
Concern with contamination Cooking raw meat Repeatedly checking meat doneness
Being around sick people Body scanning for symptoms
P4 Things she does should end in even numbers No specific triggers Counting steps when walking,
Counting syllables when talking
Hair should look even on both sides Looking at herself in the mirror Pulling hair/eye lash so that both sides look even
P5 Fear of making mistakes When worrying about something or feeling lost Rumination about how she did something wrong that directly or indirectly caused the outcome
Tightening hands to the extent that it hinders other activity
P6 Concern with contamination Accidentally touching the bathroom sink when washing hands Washing hands
Knowing some man used her bathroom Cleaning everything in the bathroom she suspects the man touched
Hearing the housekeeper washing utensils Checking if housekeeper’s hands are clean
Concern with missing important information in conversations No specific triggers Repeatedly ask the interlocutor what they just said
P7 Concern with contamination Seeing dust or stain in the kitchen Cleaning the entire kitchen
Riding public transportation Changing clothes when back home
Fear of hitting a pedestrian while driving without knowing Driving through a pothole Checking for marks on the car
Driving under low visibility on road
P8 Concern with contamination Touching things that multiple people have touched (e.g., door knob) Washing hands
Not using hands as much as possible
Fear of being responsible for something terrible happening Leaving home Putting hands under the faucet to make sure no water dripping
Hearing conversations about accidents Going back and checking if the door is locked
P9 Fear of being responsible for something terrible happening When finishing cooking Checking if the burners are off
Fear of being late Before bed time Repeatedly checking if the alarms are on
Fear of hitting a pedestrian while driving without knowing Driving down a residential street Checking rear mirror
P10 Things should be put in specific and positions No specific triggers Making sure nothing on the table is moved
Seeing baking powder on countertop Cleaning the countertop when someone is still cooking
Fear of deviating from routine would cause something terrible to happen Anything delaying his original schedule (e.g., microwave occupied by others) Counting delayed time to make up for it at the end of the day
Table A.1. Detailed personal experience of OCD (obsessions, triggers and compulsions) of reported by participants in Study 1. Only representative triggers are listed in cases where participants reported more than three triggers.
Themes Sub-themes Example codes
Trigger characterization Spectrum of triggers sound of utensil bumping, before bed time, doubt
Sensory attributes color, sticky texture, liquid more gross than solid
Affordances door knob, bar in bus, trash can with a lid
Spatial attributes proximity to trigger, amount of triggers, position/pose of trigger
Certainty of contamination origins unknown food source, who touched the food, people they trust
Other factors affecting trigger intensity mental stress, self-care, physical stress
Connecting triggers to compulsions Direct trigger interaction jiggling the door lock, moving stuff from heating vents
Consequence mitigation cleaning, changing outfit, body scanning for symptoms
Assisted compulsions seeking reassurance, asking whether telling truth
Effect of environment familiarity on compulsions more intense with closed people, compulsion less prominent in public
OCD symptom management strategies Exposure practice in daily life contaminate with battery, maintaining eye contact with triggers
Accepting & challenging obsessive thoughts uncertainty mindset, thought diffusion, focusing on values
Reducing the urge for compulsions self reminding compulsion is unnecessary, compulsion not working from experience, make compulsion more difficult
Reassurance logging taking picture, changing to 24hr time
Relaxation breathing skills, muscle relaxation
Self-compassion reflections from success, celebrate breakthrough day
External mental health support not judging shameful behaviors, family accommodating to OCD
Challenges in OCD symptom management OCD-induced social barriers OCD interrupt social interactions, downplay to get trust
Inconvenient symptom logging & tracking no self-tracking, digital journal on symptom tracking
Table A.2. Themes, Sub-themes and Example Codes generated from Study 1
Themes Sub-themes Example codes
OCD diagnosis and assessment Identify triggers interview, self-monitor triggering events
Identify compulsion untypical compulsion, mental compulsion, duration on compulsion
Psychoeducation to gain insights Psychoeducation as standalone session explain how therapy works, explain what to expect
Improve behavior awareness by homework self-monitor intrusive thoughts, self-monitor compulsion
Active treatment via ERP Building exposure hierarchy using subjective experience SUDS, ask extent to stay in anxiety, ask extent to conduct rituals
Examine trigger properties identify fear together, fluid process, identify wiggle room
Designing diverse exposures in vivo exposures, imaginal exposures, material exposures
Exposure Practices distress detection, ask mental compulsion, practice until decreased SUDS
Issues with ERP only focus on behavioral change, mental compulsion to compensate behavioral
Making long-term progress Using relaxation techniques suitably only after exposure, only for long exposure, relaxation causes hopelessness
Promoting out-of-session success via homework people not good at tracking , identify homework barriers, afraid to do homework
Technology recommendation for self-symptom management OCDefeat, low motivation, mindfulness app, concern about privacy
Table A.3. Themes, Sub-themes and Example Codes generated from Study 2

### v2

@@
-”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder
+“It was Mentally Painful to Try and Stop”: Design Opportunities for Just-in-Time Interventions for People with Obsessive-Compulsive Disorder in the Real World
 Abstract.
-Obsessive-compulsive disorder (OCD) is a mental health condition significantly affecting people’s quality of life.
-Although OCD can be effectively treated with evidence-based therapy (e.g., exposure and response prevention), Managing OCD symptoms independently, as an indispensable part of successful treatment, remains challenging due to fear confrontation and lack of appropriate support.
-We aim to comprehensively understand the challenges and needs in OCD self-management from the perspectives of both people with OCD and OCD therapists.
-Through interviews with 10 participants with diverse OCD conditions and seven therapists, we characterized different OCD symptoms, typical triggering factors, strategies, technology use, and barriers both inside and outside of therapy.
-Our findings highlighted gaps between OCD self-management needs and currently available support.
-Building on these insights, we suggest in-situ self-management technologies for OCD, including personalized symptom tracking, in-situ interventions, and discuss how OCD-specific privacy and social needs can be fulfilled with technology and beyond.
-mental health, OCD, Obsessive-Compulsive Disorder, Exposure and Response Prevention, Acceptance and Commitment Therapy, in-situ intervention
-copyright: noneccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
+Obsessive-compulsive disorder (OCD) is a mental health condition that significantly impacts people’s quality of life.
+While evidence-based therapies such as exposure and response prevention (ERP) can be effective, managing OCD symptoms in everyday life—an essential part of treatment and independent living—remains challenging due to fear confrontation and lack of appropriate support.
+To better understand the challenges and needs in OCD self-management, we conducted interviews with 10 participants with diverse OCD conditions and seven therapists specializing in OCD treatment.
+Through these interviews, we explored the characteristics of participants’ triggers and how they shaped their compulsions, and uncovered key coping strategies across different stages of OCD episodes.
+Our findings highlight critical gaps between OCD self-management needs and currently available support.
+Building on these insights, we propose design opportunities for just-in-time self-management technologies for OCD, including personalized symptom tracking, just-in-time interventions, and support for OCD-specific privacy and social needs—through technology and beyond.
+mental health, OCD, Obsessive-Compulsive Disorder, Exposure and Response Prevention, Acceptance and Commitment Therapy, just-in-time intervention
+journalyear: 2025copyright: ccconference: The 27th International ACM SIGACCESS Conference on Computers and Accessibility; October 26–29, 2025; Denver, CO, USAbooktitle: The 27th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’25), October 26–29, 2025, Denver, CO, USAdoi: 10.1145/3663547.3746394isbn: 979-8-4007-0676-9/2025/10ccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
 1. Introduction
 Obsessive-Compulsive Disorder (OCD) is a mental health disorder affecting approximately 2% of world population (Sasson et al., 1997).
-People with OCD experience uncontrollable, recurring thoughts (obsessions) usually triggered by specific events (triggers), and engage in repetitive behaviors (compulsions) to relief the anxiety caused by obsessions (of Mental Health, NIMH).
-Contrary to the common misconception that OCD is simply ’being clean’ or ’focused on order’ (Spencer and Carel, 2021), OCD symptoms are highly diverse (Hirschtritt et al., 2017) and can significantly affect quality of life especially when compulsions become severe enough to disrupt daily activities (Macy et al., 2013; Eisen et al., 2006; Subramaniam et al., 2013).
-Coping with OCD symptoms independently in daily life is crucial but challenging as real-life triggers can be too overwhelming to confront during intense OCD episodes without suitable guidance (Lind et al., 2013; Hezel and Simpson, 2019; Wheaton and Chen, 2021; Ojalehto et al., 2020).
-Many individuals with OCD seek professional support through cognitive behavioral therapy (CBT), a type of psychotherapy that help people identify and change unhelpful thoughts and behavior patterns (Beck, 2011).
-Despite its effectiveness (Yule and Whittal, 2017; Blakey et al., 2017; Hirschtritt et al., 2017), over 50% of people with OCD lack access to the therapy treatment (García-Soriano et al., 2014; Ruscio et al., 2010; Chong et al., 2012; Mayerovitch et al., 2003).
+People with OCD experience uncontrollable, recurring thoughts (obsessions) usually triggered by specific events (triggers), and engage in repetitive behaviors (compulsions) to relieve the anxiety caused by obsessions (of Mental Health, NIMH).
+Contrary to the common misconception that OCD is simply ‘being clean’ or ‘focused on order’ (Spencer and Carel, 2021), OCD symptoms are highly diverse (Hirschtritt et al., 2017) and can significantly affect quality of life especially when compulsions become severe enough to disrupt daily activities (Macy et al., 2013; Eisen et al., 2006; Subramaniam et al., 2013).
+Coping with OCD symptoms independently in daily life is crucial but challenging as real-life triggers can be too overwhelming to confront without suitable guidance (Lind et al., 2013; Hezel and Simpson, 2019; Wheaton and Chen, 2021; Ojalehto et al., 2020).
+Many people with OCD seek professional support through cognitive behavioral therapy (CBT) to receive help in identifying and changing obsessive thoughts and behavior patterns (Beck, 2011).
+Despite its effectiveness (Yule and Whittal, 2017; Blakey et al., 2017; Hirschtritt et al., 2017), over 50% of people with OCD lack access to the therapy treatment (Garc´ıa-Soriano et al., 2014; Ruscio et al., 2010; Chong et al., 2012; Mayerovitch et al., 2003).
 Moreover, many who complete the treatment continue to experience residual symptoms and struggle to apply the skills learned in therapy to their daily lives (Eddy et al., 2004).
-These challenges highlight the needs for effective self-management technology that supports people with OCD in-situ beyond traditional therapy.
-Recent advances in AI and sensing techniques have shown great potential for in-situ self-management for OCD—through intelligent scene understanding and behavior recognition (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024; Liu et al., 2024), such technology can detect triggering events and compulsive behaviors, enabling timely, in-situ interventions to help people with OCD address obsessions and resist compulsions during OCD flare-ups (Burchard et al., 2022; Olbrich et al., 2016).
-Despite the potential, research on in-situ intervention technologies for OCD remains limited and primarily focuses on specific symptoms, including recognizing excessive hand-washing on a smartwatch (Burchard et al., 2022) and detecting excessive checking behaviors via a short moving distance over an extended period of time based on GPS location (Olbrich et al., 2016).
-Given the complex and heterogeneous nature of OCD (Hirschtritt et al., 2017), timely research is needed to deeply understand the diverse personal experiences of people with OCD, uncover their challenges and needs, and identify suitable intervention feedback, thus inspiring effective, generalizable technology design.
-To fill this gap, we contribute the first comprehensive study with both people with OCD and OCD professionals.
-We first interviewed 10 participants with diverse OCD symptoms about their detailed experiences of OCD, symptom self-management strategies, persisting challenges, and desired technology to assist with self-management.
-To complement these personal insights, we further conducted an interview with seven mental health care providers who specialized in OCD to learn about their treatment strategies throughout different stages of OCD therapy and technology recommendations to facilitate symptom management outside of therapy context, guiding the feedback design for in-situ intervention technology for OCD self-management.
-Our studies detailed OCD people’s individualized triggers and compulsions and identified typical properties that predominantly affect trigger intensity.
-We found consistency between self-management strategies used by people with OCD and those taught by OCD therapists, while also uncovering nuances in treatment strategies and use of technology.
-By combining insights from both studies, we highlight gaps in OCD self-management despite currently available resources.
-By triangulating the findings from both studies, we derive technology design implications from trigger and compulsion detection, in-situ feedback design, and privacy perspectives.
-Beyond in-situ self-management, some implications can also benefit OCD therapy, such as supporting out-of-session practice.
-In summary, our research contributes: (1) the first study that thoroughly understand in-situ self-management strategies and needs from perspectives of both people with OCD and OCD professionals (2) rich qualitative data that reveals detailed personal experiences of people with OCD to inspire in-situ technology design (3) design implications that focused on input (OCD trigger and compulsion recognition) and output (feedback design) of in-situ intervention technology and OCD-specific privacy and social needs.
+These challenges highlight the needs for effective just-in-time self-management technology that supports people with OCD in real life beyond traditional therapy.
+Recent advances in AI and sensing techniques have shown great potential for just-in-time OCD interventions.
+Through intelligent scene understanding and behavior recognition (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024; Liu et al., 2024), such technology can potentially detect triggering events and compulsive behaviors and generate timely, in-situ interventions to help people with OCD address obsessions and resist compulsions during OCD flare-ups (Burchard et al., 2022; Olbrich et al., 2016).
+Despite the potential, HCI research on just-in-time intervention for OCD remains limited and primarily focuses on specific symptoms, for example, recognizing excessive checking behaviors when a short moving distance over an extended period of time is detected based on GPS location and providing audio alarm to remind the users (Olbrich et al., 2016).
+However, OCD is way more complex and heterogeneous with diverse triggers and symptoms (Hirschtritt et al., 2017).
+Timely research is needed to deeply understand the diverse triggering experiences of people with OCD, uncover their challenges and needs, and identify suitable intervention feedback, thus inspiring effective, generalizable technology design.
+To fill this gap, we contribute the first interview study with ten people with OCD and seven OCD professionals to understand the diverse triggers and experiences of OCD, the persisting challenges faced by OCD individuals, real-life coping strategies, and desired (or recommended) technology for just-in-time interventions.
+Our study cover the perspectives of both stakeholders to identify the agreement and tension between people with OCD and the professionals, thus deriving most suitable intervention design guidelines that not only fulfill user needs but also align with professional practices.
+Our study revealed unique findings towards OCD users’ experiences and preferences, as well as their tension with the professionals.
+Specifically, we found that instead of the trigger type in itself (e.g., a trash can), it was the different properties of triggers (e.g., fullness of the trashcan, whether there is a lid) that predominantly affect user experience.
+Moreover, besides certain consistency, we identified conflicts between the coping strategies adopted by people with OCD and those taught by OCD therapists.
+For example, the relaxing techniques (e.g., breathing practice) commonly used by OCD users in an OCD episode were not always recommended by the professionals.
+Our research highlights the gaps in OCD self-management despite currently available resources and derives technology design implications from trigger and compulsion detection, intervention feedback design, and ethics and privacy perspectives.
+Beyond just-in-time interventions in real life, some of our implications can also benefit OCD therapy, such as supporting out-of-session practice.
 2. Related Work
-2.1. Obsessive-Compulsive Disorder & Therapy Treatments
-2.1.1. What is OCD
+2.1. Background of OCD
 Obsessive-Compulsive Disorder occurs when a person is caught in a loop of obsessions and compulsions (American Psychiatric Association et al., 2013).
@@
 In response to these obsessions, the compulsive behaviors can be excessive hand washing, repeated checking (e.g., whether the oven is off), and repeated neutralizing thoughts (e.g., “I’m not a bad person”), respectively (Hirschtritt et al., 2017).
-Unlike obsessions and compulsions, the diversity of triggers has been overlooked in OCD research.
+However, unlike obsessions and compulsions, the diversity of triggers that cause obsessions and compulsions has been overlooked in OCD research.
 Existing studies focus on generic triggers for specific symptoms (Mataix-Cols et al., 2009) or describe OCD episodes without clearly distinguishing triggers (Knapton, 2016).
-Since triggers are highly personalized and important in characterizing OCD symptoms (Baioui et al., 2013), it is crucial to further understand and categorize triggers by exploring OCD people’s personal experiences.
-2.1.2. Common Therapy Treatments for OCD
-To combat OCD, cognitive behavioral therapy (CBT) is the most effective method (Hirschtritt et al., 2017), focusing on identifying and changing patients’ unwanted thinking and behavioral patterns (Beck, 2011).
-CBT includes different forms, such as exposure and response prevention (ERP) therapy, cognitive therapy (CT), and acceptance and commitment therapy (ACT) (Björgvinsson et al., 2007).
-ERP therapy has been widely used to facilitate behavioral changes (Blakey et al., 2017; Björgvinsson et al., 2007; Hirschtritt et al., 2017).
-It is a psychotherapy that gradually exposes individuals with OCD to triggers that provoke obsessions while simultaneously discouraging compulsions (Blakey et al., 2017; Hirschtritt et al., 2017).
-To decide the exposure tasks, an exposure hierarchy is built to cover the clients’ OCD triggers ranked based on the intensity of distress they cause.
-In between sessions, clients are also expected to complete exposure practices to achieve better outcome.
-Once the client completes the most difficult exposure successfully, the treatment can be concluded.
-However, ERP’s nature of ‘facing the fears’ can distress the clients and be time consuming (Patel et al., 2021), hence resulting in high dropout rate (Ong et al., 2016).
-To facilitate cognitive change, cognitive therapy (CT) is conducted to correct unhelpful and dysfunctional beliefs that lead to obsessions.
-Strategies are commonly developed from the underlying characteristics of the dysfunctional beliefs, such as overestimation of threat.
-An example strategy can be showing the client that the catastrophic outcome is a consequence of many steps so its probability is much lower than they expect (Yule and Whittal, 2017).
-Prior research shows that CT can achieve comparable effectiveness to ERP (Yule and Whittal, 2017), and that CT is better tolerated than ERP with lower dropout rate (Yule and Whittal, 2017; Björgvinsson et al., 2007).
-More recently, acceptance and commitment therapy (ACT), as a newly developed form of CBT, has been applied to the treatment of OCD and can significantly reduce OCD symptoms (Philip and Cherian, 2021; Evey and Steinman, 2023).
-Unlike traditional CBT (e.g., ERP), ACT encourages people to accept the presence of obsessions and the distress resulting from them, without attempting to change these feelings through compulsive behaviors (Smith et al., 2017).
-It aims to improve the ability of people with OCD to stay at the present, not attending to previous or future experiences (Smith et al., 2017; Philip and Cherian, 2021; Evey and Steinman, 2023).
-In ACT, people with OCD are guided to identify the values that are more important to them and encouraged to commit to behaviors that are consistent to these values.
-For example, a person with OCD who has identified spending time with family as a value may find excessive hand washing interfering this important event of life.
-Research has shown that combining ERP and ACT can further improve people’s ability to respond to more intense obsessions (Ong et al., 2020).
-In summary, CBT has proven to be effective in addressing OCD symptoms from both obsession and compulsion aspects.
+Since triggers are highly personalized and important in characterizing OCD symptoms (Baioui et al., 2013), it is crucial to understand and categorize triggers by exploring OCD people’s personal experiences.
+To combat OCD, cognitive behavioral therapy (CBT) is the most effective method (Hirschtritt et al., 2017), focusing on identifying and changing people’s unwanted thinking and behavioral patterns (Beck, 2011).
+CBT for OCD includes different forms, such as exposure and response prevention (ERP) therapy, cognitive therapy (CT), and acceptance and commitment therapy (ACT) (Björgvinsson et al., 2007).
+ERP helps people with OCD gradually face their OCD triggers while refraining from compulsions, aiming for behavioral change through repeated exposure to those triggers (Blakey et al., 2017; Björgvinsson et al., 2007; Hirschtritt et al., 2017).
+Cognitive Therapy (CT) focuses on identifying and restructuring dysfunctional beliefs that lead to obsessions, helping individuals reevaluate perceived catastrophic outcome (Yule and Whittal, 2017).
+Acceptance and Commitment Therapy (ACT), in contrast, encourages acceptance of obsessive thoughts without engaging in compulsions and promotes value-driven behavior to reduce the impact of OCD symptoms (Smith et al., 2017, 2017; Philip and Cherian, 2021; Evey and Steinman, 2023).
+CBT has proven to be effective in addressing OCD symptoms from both obsession and compulsion aspects.
 However, it is not available 24/7 whenever people need support to manage their issues (Blakey et al., 2017).
-The homework between sessions (e.g., symptom self-monitoring, self-guided exposure practices)—as an important part of CBT—could be difficult for clients to comply, affecting the outcome of the therapy (Wheaton and Chen, 2021).
+The homework between sessions (e.g., symptom self-monitoring, self-guided exposure practices)—as an important part of CBT—could be difficult for clients to comply, affecting the outcome of therapy (Wheaton and Chen, 2021).
+Despite the drawbacks, the strategies used in these therapies can potentially bring insights for just-in-time interventions.
+Our research thus seize this opportunity to conduct studies with people with OCD to understand their experiences and coping strategies, triangulated with professionals’ perspectives, to inspire real-world intervention technologies beyond conventional therapy.
 2.2. Technology for OCD Symptom Self-Management
-To overcome the challenges of OCD symptom self-management, a variety of different forms of new technology have been developed and evaluated.
-Existing efforts have been focused on symptom and behavior monitoring(Khan et al., 2011; Frank et al., 2023), proactive skill building(Khan et al., 2011; Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010), and in-situ intervention (Burchard et al., 2022; Olbrich et al., 2016).
-We discuss prior work on self-symptom management following this taxonomy.
+To overcome the challenges caused by OCD in real life, different forms of technologies have been developed and evaluated.
+Existing efforts have mainly focused on symptom and behavior monitoring (Khan et al., 2011; Frank et al., 2023) and proactive skill building (Khan et al., 2011; Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010), but limited work has focused on just-in-time interventions for specific OCD symptoms (Burchard et al., 2022; Olbrich et al., 2016).
+We discuss prior work on OCD self-management technology below.
 2.2.1. Symptom monitoring
-Symptom monitoring is an effective approach to supporting self-management for various mental health conditions, such as anxiety disorder and depression (Kruzan et al., 2023), due to its potential of improving a person’s symptom (e.g., thoughts, behavior) awareness (Kruzan et al., 2023).
-In the context of OCD, prior work has explored two ways of self-monitoring, (1) self-report, and (2) passive sensing.
-Ecological Momentary Assessment (EMA) — a research approach in clinical psychology to periodically prompting for people’s behaviors and experiences in their natural environments — has been broadly used to track and assess an OCD person’s symptoms (Frank et al., 2023; Brown et al., 2020; Tilley and Rees, 2014; Rupp et al., 2020), and demonstrated high acceptability, practicability, and representativeness (Frank et al., 2023).
-However, research has shown that overly relying on self-report approach can result in noisy (Frank et al., 2023) and missing data.
+Symptom monitoring is an effective approach to supporting mental health self-management, such as anxiety disorder and depression (Kruzan et al., 2023), due to its potential improvement in symptom (e.g., thoughts, behavior) awareness (Kruzan et al., 2023).
+Specifically in the context of OCD, prior work has explored two ways of self-monitoring, (1) self-report, and (2) passive sensing.
+The most common self-report approach is via Ecological Momentary Assessment (EMA), which periodically prompts the users to track and assess their own symptoms in their natural environments (Frank et al., 2023; Brown et al., 2020; Tilley and Rees, 2014; Rupp et al., 2020).
+While this method has demonstrated high acceptability, practicability, and representativeness (Frank et al., 2023), research has shown that overly relying on self-report approach can result in noisy and missing data (Frank et al., 2023).
 In contrast, passive sensing enables rich data collection without burdening users.
-Researchers have investigated passive symptom monitoring using various sensors (Thierfelder et al., 2022; Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022).
+Researchers have investigated passive symptom monitoring using various sensors (Thierfelder et al., 2022; Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022; Lønfeldt et al., 2023).
 For example, Thierfelder et al.
-(Thierfelder et al., 2022) monitored OCD-induced stress by measuring heart rate, hand activity and gaze fixation, and demonstrated that stress and anxiety detection is possible using multimodal sensors.
-Other studies measured sleep quality of people with OCD using actigraphy (Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022).
-While symptom monitoring can improve symptom awareness, prior work has primarily focused on monitoring users’ internal states instead of contextual information which is important in characterizing OCD symptoms (Llorens-Aguilar et al., 2021).
-In addition, without proper knowledge and coping skills for symptom management, as well as actionable insights derived from collected data, the effectiveness of symptom monitoring can be further limited.
+(Thierfelder et al., 2022) monitored OCD-induced stress by measuring heart rate, hand activity and gaze fixation, and demonstrated the possibility of using multimodal sensory data to detect stress and anxiety.
+Other studies have focused on monitoring different aspects of OCD symptoms, including compulsion (e.g., excessive hand-washing) detection using inertial motion sensor data on a smartwatch (Wahl et al., 2022; Burchard et al., 2022), OCD episode tracking using blood volume pulse and skin temperature (Lønfeldt et al., 2023; Olesen et al., 2023), and sleep quality measurement using actigraphy (Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022).
+While symptom monitoring can improve symptom awareness, prior work has primarily focused on monitoring users’ internal states instead of contextual information (e.g., potential triggers in the environment), which is important in characterizing OCD symptoms (Llorens-Aguilar et al., 2021).
+In addition, without proper knowledge and coping skills for symptom management, the effectiveness of symptom monitoring can be further limited.
 2.2.2. Proactive Skill Building
-To support out-of-session symptom management skill building, smartphone-based intervention has been developed that incorporate evidence-based therapy treatment in peoples’ everyday life (Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010).
-LiveOCDFree is an app that guides the user to set up and conduct self-administered ERP by creating an exposure hierarchy, and identifying specific exposures to practice (e.g.
-touching doorknobs).
+To support skill building outside therapy sessions, virtual therapy applications based on desktop or smartphone have been developed (Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010).
+For example, LiveOCDFree is an app that guides the user to set up and conduct self-administered ERP therapy by creating an exposure hierarchy and identifying specific triggers to practice (e.g., touching doorknobs).
 The user is able to set reminders for ERP practice and monitor their progress.
 An evaluation with 21 participants demonstrated significant improvement in OCD and anxiety symptoms after using the app (Boisseau et al., 2017).
-Similarly, NOCD supported self-administered ERP between therapy sessions incorporating ACT strategies.
-A study with 2,069 participants showed that NOCD effectively reduced OCD symptoms (Feusner et al., 2021).
-However, there are limitations with self-administered ERP such as inflexible template for specifying obsessions and compulsions which restricts user autonomy (Hong et al., 2020).
-Prior work has also reported clinical significance of web-based CBT in reducing OCD symptoms (Ferreri et al., 2019), such as BT Steps, a program that guides users with OCD to learn symptom management skills (Schure et al., 2020).
-These technologies effectively supported OCD people’s symptom management outside of therapy.
-However, they all require OCD clients to proactively practice to build up coping skills, instead of providing targeted support for clients to respond to intense real life triggers, which can be difficult to manage without therapists’ help (Lind et al., 2013).
-In contrast, in-situ intervention technologies that implement therapeutic strategies (e.g., exposure practice) along with real life events (Kim et al., 2024) can potentially assist with OCD treatment in a more effective manner.
-2.2.3. In-Situ Intervention
+Similarly, NOCD supported self-administered ERP between therapy sessions by incorporating ACT strategies.
+A study with 2,069 participants showed that it effectively reduced OCD symptoms (Feusner et al., 2021).
+Beyond conventional applications, recent advances in large language models (LLMs) further presents new opportunities in delivering personalized mental well-being support (Guo et al., 2024; Kian et al., 2024; Nie et al., 2024; Iftikhar et al., 2024).
+For example, Kian et al.
+(Kian et al., 2024) developed an LLM-powered assistive robot to deliver out-of-session CBT exercises for users with depression and achieved comparable effectiveness to traditional methods.
+All above therapy simulation technologies require users to proactively practice to build up coping skills.
+However, triggers in real world could be much more unpredictable and intense, which can still be hard to manage even with certain skills established (Eddy et al., 2004).
+In contrast, targeted, just-in-time interventions that are adaptive to real world events (Kim et al., 2024) can potentially be more effective in real life.
+2.2.3. Just-in-time Interventions
 As opposed to pre-scheduled therapy sessions, the increasingly powerful mobile and sensing technologies allow for more flexible and intelligent mental well-being interventions.
-Such systems can recognize and adapt to users’ internal mental states and physical context in daily life to provide just-in-time mental support — an intervention design that provides the right type and amount of support, at the right time, in the context that the person needs it most and is most likely to be receptive(Nahum-Shani et al., 2018).
-Such interventions have been applied to treat different mental health conditions, such as substance use disorder (Gustafson et al., 2014), depression (Burns et al., 2011), and schizophrenia (Ben-Zeev et al., 2014, 2013).
-For example, ACHESS (Gustafson et al., 2014) is a mobile application that tracks the GPS location of a user with alcohol use disorders and provides text-based support when they go to high-risk locations (e.g., a bar).
+Such systems can recognize and adapt to users’ internal mental states and physical context in daily life and provide just-in-time support—an intervention design that delivers the right type and amount of support, adaptive to the timing and context when the user needs it.
+(Nahum-Shani et al., 2018).
+Just-in-time interventions have been applied to various mental health conditions, such as substance use disorder (Gustafson et al., 2014), depression (Burns et al., 2011; He et al., 2024; Zhao et al., 2023; Costa et al., 2016), schizophrenia (Ben-Zeev et al., 2014, 2013), and excessive phone use (Xu et al., 2022) For example, ACHESS (Gustafson et al., 2014) is a mobile app that tracks the GPS location of a user with alcohol use disorders and provides text-based reminder when they go to high-risk locations (e.g., a bar).
 An evaluation with 349 participants shows that ACHESS significantly reduced risky drinking days.
-There has been limited exploration on in-situ intervention for OCD.
-Burchard et al.
-(Burchard et al., 2022) develops a deep-learning-based system to detect excessive hand-washing—a typical compulsion in contamination OCD—using the inertial motion sensor data on a smartwatch.
-However, although the system supports real-time symptom recognition, it does not provide in-situ feedback to the user.
-Olbrich et al.
-(Olbrich et al., 2016) develops a smartphone application Geo-Feedback that actively detects the user’s excessive checking behavior indicated by short moving distance over an extended period of time via GPS location and evokes a selected alarm when the excessive checking behavior is detected.
-A case study with a person with severe OCD indicates the effectiveness of the application.
-These examples have demonstrated initial success in designing real-time interventions for people with OCD.
-However, they are purely focused on physical compulsions and highly limited in context, hindering generalization to other OCD conditions.
-More research is needed to guide the design of in-situ intervention technology, including both the detection of OCD triggers and the design of effective and acceptable interventions consistent with therapeutic advice.
-3. Study 1: Interviewing People with OCD
-We first conducted semi-structured interviews with 10 participants with OCD.
-We aimed to comprehensively understand the impact of OCD (e.g., obsessions, compulsions, triggers) on their daily life, with an ultimate goal of distilling the needs and challenges of OCD symptom self-management and informing the design of in-situ intervention.
-The study was approved by Institutional Review Board (IRB) at our university.
-3.1. Method
-3.1.1. Participants
-We recruited 10 participants (7 female and 3 male) whose ages ranged from 19 to 47 ($Mean$ = 28.5, $SD$ = 8.5).
-Participants’ demographic information and OCD diagnosis are detailed in Table 1.
-We recruited participants from online OCD support groups and the research email service at our university.
-Participants were screened via an email to (1) inform them of the interview content to ensure they would be comfortable participating in this study, and (2) check their eligibility.
-A participant was eligible if they were (1) at least 18 years old, and (2) diagnosed with mild to moderate OCD and were willing to disclose their OCD symptoms with us.
-We did not recruit people experiencing severe OCD at the time of the study, because our interview questions can be triggering without adequate ability to manage symptoms independently.
-Participants were compensated for \$20 per hour.
-At the time of the study, all participants except P6 (from India) were located in the United States.
+Zhao et al.
+(Zhao et al., 2023) developed a wearable device that effectively reduced anxiety during high-stress situations through affective touch—a tactile intervention that renders a soft stroking sensation on the user’s forearm.
+Moreover, Xu et al.
+(Xu et al., 2022) designed TypeOut, which integrates value-based self-affirmation messages with a typing-based app unlock process, significantly reducing app usage and screen time.
+Despite the exploration of just-in-time interventions for other mental health conditions, the work for OCD remains nascent.
+The only work we know is from Olbrich et al.
+(Olbrich et al., 2016), who develops a smartphone app Geo-Feedback that actively detects users’ excessive checking behaviors indicated by short moving distance over an extended period of time via GPS location and evokes an alarm when such excessive behaviors are detected.
+A case study with one user with severe OCD indicates the effectiveness of the app.
+While demonstrating the initial success of just-in-time interventions for people with OCD, this work focuses only on one specific OCD condition.
+More research is needed to understand the diverse, complex nature of OCD as well as user needs, thus deriving more generalizable guidelines for just-in-time intervention technology (including both trigger detection and intervention design) to support various OCD types in real life.
+3. Method
+To better understand how OCD (e.g., obsessions, compulsions, triggers) impacts people’s daily life and to identify effective just-in-time intervention strategies, we conducted in-depth interviews with 10 people with OCD (P1-P10) and seven OCD therapists (T1-T7).
+Our goal was to distill the triggering events, strategies, and challenges of OCD self-management in real life, ultimately informing the design of just-in-time intervention technologies.
 ID Age/ Gender Severity Diagnosed Time OCD Types
@@
 P10 35/M Moderate 6 months ago Just-right OCD
-Table 1. Demographic information and basic obsession types of the 10 participants living with OCD in Study 1.
-Our participants demonstrated diverse OCD types (Table 1), covering most common obsessions and compulsions in the Y-BOCS Symptom Checklist (Goodman et al., 1989).
-Six (P1, P5, P6-9) reported to have responsibility OCD, whose obsessions were around not being responsible enough or fear of making mistakes.
-Common compulsions reported by participants included checking if the door is properly locked.
-Five participants (P1, P3, P6-P8) had contamination OCD (e.g., dirt, germ) with common compulsions being washing hands and cleaning.
-P4 and P10 had just-right OCD which manifested as being obsessed with symmetry or the position of objects.
-The corresponding compulsion would be fixing asymmetry (P4) and objects that are not where they should be (P10).
-P2 had harm OCD, having intrusive thoughts about harm occurring to herself and people around her with compulsion being rumination about the negative thoughts and avoidance of the trigger.
+Table 1. Demographic information and basic obsession types of the 10 participants living with OCD.
+ID Age/ Gender Certificate Years of Experience Specialization OCD Clients Severity Level Practice Modality
+T1 46/F Licensed Clinical Social Worker 25 years OCD, anxiety disorders, eating disorders, PTSD, depression, substance use disorder Mild to severe Hybrid
+T2 47/F Clinical Psychologist 14 years OCD, axiety disorder Mostly moderate to severe Virtual
+T3 36/M License Professional Counselor 9 years OCD, anxiety disorder, phobias, health anxiety Mild to severe, mostly moderate Virtual
+T4 60/F Licensed Clinical Social Worker 13 years OCD, anxiety, depression Mostly moderate Hybrid
+T5 51/F Professional Counselor Training License (LPC-IT) 2 years and 4 months OCD, anxiety disorder, depression Mild to severe, mostly moderate Hybrid
+T6 38/M License Professional Counselor 17 years OCD, anxiety disorders, eating disorders, PTSD Mild to severe Hybrid
+T7 29/F License Professional Counselor 2 years OCD, anxiety disorder Moderate to severe Virtual
+Table 2. Demographic information and professional background of the seven OCD professionals.
+3.1. Participants
+Our study involved two key stakeholders in the OCD domain—individuals who experience OCD and professionals who specialize in OCD treatment—to develop a comprehensive understanding of the current practices and challenges, ensuring that our design insights would integrate both lived experience of people with OCD and professional recommendations.
+3.1.1. People with OCD
+We recruited 10 participants with OCD (7 female and 3 male) whose ages ranged from 19 to 47 ($Mean$ = 28.5, $SD$ = 8.5).
+Participants’ demographic information and OCD types are detailed in Table 1.
+While our participant gender coverage skewed towards women, it was consistent with the higher prevalence of OCD in female adults than male adults (Mathes et al., 2019).
+We recruited participants from online OCD support groups and the research email service at our university.
+A participant was eligible if they were (1) at least 18 years old, (2) diagnosed with mild to moderate OCD, and (3) were willing to disclose their OCD symptoms with us.
+We did not recruit people experiencing severe OCD at the time of the study because our interview questions can be triggering without adequate ability to manage symptoms independently.
+Participants were compensated for \$20 per hour.
+Our participants demonstrated a wide range of OCD types (Table 1), covering most common obsessions and compulsions in the Y-BOCS Symptom Checklist (Goodman et al., 1989).
+Six participants (P1, P5, P6-9) had responsibility OCD, whose obsessions were around not being responsible enough or fear of making mistakes; their common compulsions included checking if the door is properly locked.
+Five participants (P1, P3, P6-P8) had contamination OCD (e.g., dirt, germ) with common compulsions being excessive hand washing and cleaning.
+Two (P4, P10) had just-right OCD who were obsessed with symmetry or specific position of objects with corresponding compulsions of fixing asymmetry (P4) and objects that are not where they should be (P10).
+P2 experienced harm OCD, having intrusive thoughts about harm occurring to herself and people around her with compulsion of ruminating about the negative thoughts and avoiding the triggers.
 She also experienced sex-related obsessive thoughts (sexual OCD) when interacting with men and avoided looking at them to ease her distress.
-Table A.1 detailed the OCD experiences reported by participants.
-Our participant demographic was skewed towards women.
-However, it is consistent with the higher prevalence of OCD in adult females than adult males (Mathes et al., 2019).
-Similarly, consistent with the distribution of different OCD subtypes (Ball et al., 1996), more than half of our participants had contamination OCD and responsibility OCD.
-3.1.2. Procedure.
-The study was a single-session interview via Zoom that lasted 90 minutes.
-We first asked participants about their demographic information and diagnosis of OCD.
-Then we asked participants about their detailed personal experience with OCD.
-We prompted participants for their obsessions, triggers, and compulsions in different living environments (e.g., home, workplace, public spaces).
-For triggers, we specifically asked for finer-grained details with a goal of identifying the characteristics and rationales of triggers.
-Participants were prompted to describe strategies to manage their OCD symptoms either recommended by therapists or developed by themselves.
-We also asked about their therapy experiences and use of other resources (e.g., technologies) to assist with symptom management.
-Finally, participants discussed ideas about desired technological support for OCD symptom self-management.
-3.1.3. Data recording and analysis.
+Consistent with the distribution of different OCD subtypes (Ball et al., 1996), more than half of our participants had contamination OCD and responsibility OCD.
+Table A.1 detailed participants’ OCD subtypes.
+3.1.2. OCD Professionals
+We recruited seven OCD professionals (5 female, 2 male).
+Their ages ranged from 29 to 60 ($Mean$ = 43.9, $SD$ = 10.4).
+All participants had extensive experience working as mental health providers (at least two years), with four (T1, T2, T4, T6) having more than 10 years of experiences.
+Participants had experiences in treating OCD across various severity levels (from mild to severe) and in different practice modalities (both hybrid and virtual).
+Table 2 shows participants’ detailed information.
+We recruit OCD professionals by sending emails to local OCD therapists listed on the International OCD Foundation, a nonprofit organization providing resources and education for people with OCD (IOCDF, 2024).
+Participants were eligible if they were over 18 and had experience conducting OCD related therapy.
+We limited the recruitment to individuals who spoke English and located in the United States.
+Each participant received \$40/hour as compensation.
+3.2. Procedure
+We conducted single-session semi-structured interviews with both participants with OCD and OCD professionals via Zoom.
+Each session lasted for about 90 minutes.
+Our interview focused on different perspectives between the OCD individuals and the professionals.
+For individuals with OCD, we focused on their unique triggers, experiences, challenges, and coping strategies in daily life.
+Specifically, we first asked about their demographic information (e.g., age, gender identity) and diagnosis of OCD.
+Then we asked participants about their detailed personal experience with OCD—we prompted them about their obsessions, triggers, and compulsions in different living environments (e.g., home, workplace, public spaces).
+For triggers, we specifically asked for fine-grained details with a goal of identifying the characteristics and rationales of the triggers.
+We then asked participants about their strategies and technology use to manage their OCD symptoms in daily life, either recommended by therapists or developed by themselves.
+We also asked about their therapy experiences and use of other resources to assist with symptom management.
+Finally, participants discussed ideas about desired real-life interventions for symptom self-management.
+For OCD professionals, we focused on their strategies and practices in OCD therapy sessions to potentially inspire intervention design.
+We first asked about professionals’ demographic information, certifications, and experience as a mental health therapy provider.
+We then asked about the types of therapy they conducted, the overall therapy process (e.g., treatment stages), the strategies and technologies used at different treatment stages, and how they evaluated client progress.
+In addition to in-session practices, we asked therapists about out-of-session strategies they recommend, such as assigned homework and external support.
+We then had them discuss their current challenges in therapy practices.
+Lastly, we asked therapists about their familiarity with and opinions on emerging technologies, such as phone-based self-administered therapy.
+When therapists were unfamiliar with certain technologies (e.g., virtual reality), we provided a brief, neutral description to ensure informed responses and minimize misconceptions.
+3.3. Analysis
 We audio recorded and transcribed all interviews via an online transcription service (otter.ai [^1] https://otter.ai).
 One researcher manually cleaned all transcripts and corrected transcribing errors by checking the recorded videos.
-We analyzed the data using thematic analysis (Clarke et al., 2015; Braun and Clarke, 2006) We selected two representative transcripts as samples, and two researchers coded the samples independently with open coding.
-Then, they discussed and reconciled their codes to resolve any differences and developed an initial codebook upon agreement.
-One researcher then coded the remaining eight transcripts based on the codebook.
-If a new code emerged, the two researchers discussed on the code and updated the codebook upon agreement.
-We derived themes using a hybrid of inductive and deductive approaches (Swain, 2018; Braun and Clarke, 2006).
-First, we generated high-level themes based on fundamental concepts of OCD symptoms and management (e.g., OCD symptom management strategies) (Wu and Lewin, 2017; Knapton, 2016) using a deductive approach (Crabtree and Miller, 1992).
-Next, we used an inductive approach (Boyatzis, 1998) to generate sub-themes and identify new themes (e.g., connecting triggers to compulsions) by grouping related codes via affinity diagramming (Clarke et al., 2015; Braun and Clarke, 2006).
-Two researchers then cross-referenced the original data, codebook, and themes to make final adjustment, ensuring all codes were grouped in the correct themes.
-Our analysis resulted in over 400 codes and 23 themes and sub-themes.
-We present our themes, subthemes and example codes in Table A.2.
-3.1.4. Ethical considerations and reflexivity statement.
+We analyzed the data using thematic analysis (Clarke et al., 2015; Braun and Clarke, 2006).
+We selected two representative transcripts from OCD participants and two from OCD professionals as samples to develop initial codebooks.
+Two researchers coded the samples independently with open coding.
+Then, they discussed and reconciled their codes to resolve any differences and developed two initial codebooks upon agreement, one for participants with OCD and one for OCD professionals.
+One researcher then coded the remaining transcripts from the two groups of participants based on the codebooks, respectively.
+If a new code was constructed, the two researchers discussed on the code and updated the codebooks upon agreement.
+We first derived themes for OCD participants and OCD professionals separately, where we used a hybrid of inductive and deductive approaches, a method commonly adopted in qualitative research (Oewel et al., 2024; Swain, 2018; Braun and Clarke, 2006).
+We started with high-level themes based on fundamental concepts of OCD triggers, compulsions, therapy treatment, and self-management identified in prior work (Wu and Lewin, 2017; Knapton, 2016; Blakey et al., 2017; Wu and Lewin, 2017; Yule and Whittal, 2017) using a deductive approach (Crabtree and Miller, 1992).
+We then used an inductive approach (Boyatzis, 1998) to generate sub-themes and identify new themes by grouping related codes via affinity diagramming (Clarke et al., 2015; Braun and Clarke, 2006) for each participant group.
+Finally, we generated final themes and sub-themes across participants by comparing and combining themes from the two groups.
+Two researchers cross-referenced the original data, the two codebooks, and themes to make final adjustment, ensuring all codes were grouped in the correct themes.
+Our analysis resulted in over 700 codes and 20 themes and sub-themes.
+3.4. Ethical Considerations & Reflexivity Statement
 Our interviews involved discussing participants’ OCD symptoms, which could cause mental distress.
-To provide a safe and comfortable environment for participants, we consulted a clinical psychologist (a co-author) with experience treating OCD to ensure appropriate language was used in the interview.
-We applied people-first language (Sharif et al., 2022), such as people with OCD instead of OCD patients, to promote inclusion and reduce stigma.
-Participants were informed they could skip any uncomfortable questions and turn off their camera if preferred.
+To provide a safe and comfortable environment for OCD participants, we consulted a clinical psychologist (a co-author) with experience treating OCD to ensure appropriate language was used in the interview.
+OCD participants were informed that they could skip any uncomfortable questions and turn off their camera if preferred, and they could exit the study anytime when needed.
 We recognize that researchers’ identities and background may influence our data interpretation.
@@
 His background and lived experience with OCD enabled us to develop appropriate questioning routes, better empathize with participants, and interpret their responses via a unique lens.
-Moreover, the researcher team comprises diverse expertise, including HCI, accessibility, intelligent system design, and OCD therapy, enabling us to better contextualize the study data in existing technology and mental health literature and derive suitable technology implications to support people with OCD.
-3.2. Findings
-In this section, we first report participants’ detailed experience of OCD in their daily lives and coping strategies.
-We then unpack the challenges faced by participants despite currently available support.
-3.2.1. What Makes Something a Trigger?
-We first investigated the OCD triggers (Table A.1) to determine what to recognize and how to recognize triggers for self-management technology for OCD.
-With this goal in mind, we categorize OCD triggers into different types and examine the key attributes that contribute to a trigger, which has been overlooked in prior literature in OCD (See Section 2.1.1).
-Spectrum of Triggers.
-OCD triggers are highly idiosyncratic (Baioui et al., 2013).
-While most triggers reported by participants are specific physical objects (i.e., triggering objects), such as trash cans for contamination OCD (P1), and knives for harm OCD (P2), participants revealed other types of triggers, including contexts, sounds, and even thoughts.
+Moreover, the research team comprises diverse expertise, including HCI, accessibility, intelligent system design, and OCD therapy, enabling us to better contextualize the study data in existing technology and mental health literature, thus deriving suitable technology implications to support people with OCD.
+We interviewed both people with OCD and therapists who treat OCD to gather diverse perspectives on symptom self-management strategies and needs.
+While therapists offer insights grounded in treatment principles and their experience with clients, we acknowledge their positionality as mental health professionals rather than individuals with lived experience.
+Their interpretations may not fully capture, and may at times diverge from, how people with OCD understand and relate to their experiences.
+As such, our analysis prioritized the narratives of people with OCD and used therapist perspectives as a complementary lens.
+When tensions between the two groups emerged, we present them transparently without privileging therapists’ views.
+4. Findings
+Our study findings go beyond the conventional clinical view of OCD and takes a unique sociotechnial lens to understand the OCD individuals’ unique triggers, compulsions, and coping strategies as well as highlighting the overlaps and contrast between the perspectives of people with OCD and the professionals.
+We detail our findings below.
+4.1. Redefining OCD Triggers
+We first investigated the OCD triggers to determine what triggers to recognize and how to recognize triggers for just-in-time interventions for OCD.
+We highlight that triggers go beyond “a common list” followed by prior work (Mataix-Cols et al., 2009) but include various types in different forms.
+Moreover, it is not the trigger type (e.g., a trashcan) in itself but rather its properties (e.g., fullness of the trashcan) that can significantly affect an OCD individual’s experiences.
+4.1.1. Triggers can be anything
+Unlike prior work that limits OCD triggers to a set of ”common triggers” based on basic OCD themes (e.g., contamination, harm) (Mataix-Cols et al., 2009), our findings show that OCD triggers are highly individualized.
+While most triggers reported by OCD participants are specific physical objects (i.e., triggering objects), such as trash cans for contamination OCD (P1) or knives for harm OCD (P2), participants also revealed less recognized trigger forms, including contexts, sounds, and even thoughts.
 We found that OCD can be triggered by participants’ physical context, such as specific locations or times (P1, P3, P7-9).
@@
 Interestingly, weather can also be a trigger.
-As P7 commented, her fear of hitting a pedestrian emerged with low visibility on the road during rain or snow.
+For example, P7 indicated her fear of hitting a pedestrian emerged with low visibility on the road during rain or snow.
 OCD triggers can also be audible (P3, P6-8, P10).
-For example, coughing sound can immediately trigger P3’s contamination OCD, and accident-related keywords in a conversation would prompt P8 to check his own apartment for safety.
+For example, coughing sound can immediately trigger P3’s contamination OCD, and accident-related keywords in a conversation would trigger P8 to check his own apartment for safety.
 In addition, some participants reported mental triggers (P5), such as unspecified worries, without clear links to physical objects.
 Moreover, OCD episodes could sometimes occur even without identifiable triggers (P1, P4, P6, P10).
-Through a deeper investigation into triggers, we found that for object triggers, different properties of a trigger can significantly affect trigger intensity (i.e., the degree of distress caused by a specific trigger), rather than the object itself, leading to different levels of anxiety and compulsion.
-We summarize the following trigger properties across different types of obsessions.
+Inspiration:
+Inspiration: The diverse forms of triggers reveal the difficulty of real-time trigger recognition and highlight the needs for the use of multi-modal sensory data in trigger identification.
+4.1.2. Trigger Intensity: Properties over Objects
+Beyond the trigger type, the intensity of triggers (i.e., the mental distress induced by a certain trigger) plays an important role in determining whether an individual would be triggered.
+Our OCD professionals emphasized the importance of asking clients not just about the triggers themselves, but also about the factors that influence their intensity (T2, T3, T5).
+Specifically, the OCD professionals often work with the clients to create a trigger hierarchy—a ranked list of OCD triggers ordered based on their intensity.
+However, this process usually follows a specific template by asking about clients’ experience with a checklist of “common triggers” (e.g., Y-BOCS (Goodman et al., 1989)) and then asking them to rate their distress caused by a specific trigger (T1-T2, T4-T7) or their urge to performance compulsion (T3).
+While some professionals allowed their clients to add more diverse triggers by logging their encounters outside the therapy session (T1–T5, T7), we found that there was a lack of comprehensive method in clinic to identify all possible triggers.
+Moreover, the method of evaluating trigger intensity was completely based on client’s subject ratings, without systematic characterization of the factors that affect trigger intensity.
+Despite the lack of clinical standard, our study data from OCD individuals revealed that, for physical triggers, instead of the object type in itself, it was often the specific properties of the object that determined the trigger intensity, leading to varying levels of anxiety and compulsions.
+Below, we summarize key properties that affect trigger intensity from the experiences of different participants with OCD.
 Sensory Attributes.
-Visual attributes such as color can alter the intensity of a trigger (P3, P7).
+Participants reported multiple sensory attributes that can alter the intensity of a trigger.
+Visual attributes such as color can affect people’s perception of a trigger (P3, P7).
 For example, P7’s contamination OCD were triggered when the dirt presents high contrast against its background, such as coffee grind on a white countertop.
-While the same amount of salt would not be as triggering.
+While the same amount of white salt (also considered as dirt) with low contrast to the countertop would not be as triggering.
 In addition to color, some specific shapes can also be more triggering than others, such as sharp edges.
-P2 felt more uncomfortable when interacting with objects that have sharp tips in her home environment, including knives, pencils, and the sharp corners of newel posts on a staircase.
-Moreover, tactile attributes of a triggering object is another factor that affects trigger intensity (P1, P7).
-The sticky texture of sauce (e.g., ketchup) caused more contamination concerns to P1 and P7 than solid food.
+P2 felt more uncomfortable when interacting with objects that have sharp tips in her home environment, including knives with sharp blade, pencils with sharp tips, and the newel posts with sharp corners on a staircase.
+Moreover, tactile attributes of an object could be another factor that affected trigger intensity (P1, P7).
+For instance, the sticky texture of sauce (e.g., ketchup) caused more contamination concerns to P1 and P7 than solid food.
 Affordances.
-We found triggering obejcts that afford direct contact with the user can cause higher level of distress to participants with contamination OCD (P1, P3, P6-8).
+We found triggering objects that afforded direct contact with the user can cause higher level of distress to participants with contamination OCD (P1, P3, P6-8).
 For example, door handles in public spaces and handrails in transportation are common triggers because it is meant to be touched using hands by many people, thus considered dirtier than surfaces not requiring direct contact (e.g., automatic doors).
@@
 Spatial attributes of a triggering object is another property that determines trigger intensity.
-First, the quantity of triggering obejcts can affect the intensity of contamination OCD.
+First, the quantity of triggering objects can affect the intensity of contamination OCD.
 P3 tended to avoid occasions with big crowds of people more than others due to concerns about contagious diseases.
@@
 P2’s sexual intrusive thoughts were more likely triggered when a man was in a position that “resembled those in her intrusive thoughts.”
-Certainty of Contamination Origins.
-For participants with contamination OCD, trigger intensity was influenced by the perceived certainty of the contamination’s origin (P1, P8).
-For instance, P1 was not triggered by food residues on a plate if she knew who was responsible for the mess.
-Similarly, P8 was more concerned about germs brought by strangers to his apartment than by his own family, whom he trusted for their hygiene.
-Other Factors Affecting Trigger Intensity.
+Familiarity.
+Trigger intensity varied depending on participants’ familiarity with their triggers.
+For participants with harm OCD (P2), familiar triggers can be more intense.
+For example, P2 experienced more intrusive thoughts about accidentally hurting people when around family or friends rather than strangers.
+Conversely, participants with contamination OCD (P1, P8) found familiar triggers less intense.
+For instance, P1 was not distressed by food residue if she knew who was responsible for it.
+“Right now, in my kitchen, there’s a few dirty dishes in the sink, and there is a pot on the stove, and a box of pasta that needs to be put away, but I know that my roommate made Pasta last night, and so I’m not at all stressed.
+Because I know exactly where everything came from.” Similarly, P8 was more concerned about germs brought into his apartment by strangers than by his own family, whose hygiene he trusted.
+Internal Factors Affecting Trigger Intensity.
+In addition to trigger properties, participants with OCD highlighted several internal factors that influenced trigger intensity.
 Six participants (P1-4, P6, P7) further pointed out that their OCD triggers intensified when experiencing mental stress.
-It can be from schoolwork (P1, P2, P4), unfamiliar environment (P1) and personal life (P4).
-Additionally, poor self-care such as sleep deprivation and physical inactivity is another factor that exacerbates trigger intensity (P1, P2).
+It can be from schoolwork (P1, P2, P4), being in unfamiliar environments (P1) and personal life issues (P4).
+Additionally, poor self-care, such as sleep deprivation and physical inactivity, also exacerbated trigger intensity (P1, P2).
 Interestingly, unlike mental stress, physical stress could sometimes help reduce trigger intensity (P7, P8).
 As P8 explained, “if I’m pretty tired, my brain can’t put so much energy towards the OCD.” Similarly, trigger intensity can be reduced when short-term memory is fully occupied by other tasks (P2, P6), e.g., when working (P6).
-In summary, trigger intensity is affected by not only trigger itself, but also properties of triggers whose variation can result in significantly different anxiety levels.
-Since OCD symptoms are highly personalized (Hirschtritt et al., 2017), trigger properties should be accounted for when designing systems for trigger recognition.
-Moreover, given the diverse types of triggers beyond physical objects (e.g., sounds, contexts), multi-sensing techniques and multi-modal recognition methods should also be considered in trigger detection.
-3.2.2. Connecting Triggers to Compulsions
-Besides triggers, compulsion is another important indicator to trigger in-situ interventions for self-management technology.
-Unlike prior OCD literature that merely focused compulsions on user behaviors (e.g., washing hands) (Starcevic et al., 2011; Goodman et al., 1989), we found that our participants constantly associated their compulsion behaviors with triggers.
-We thus categorize compulsions into five categories based on their relationship with triggers to inform compulsion recognition for in-situ intervention technology, including: 1) direct trigger interaction, 2) consequence mitigation, 3) trigger avoidance, 4) assisted compulsions, and 5) mental compulsions.
-As trigger avoidance (e.g., avoiding contact with objects deemed dirty by people with contamination OCD) and mental compulsion (e.g., rumination) has been recognized in existing OCD research (Association, 2024; Williams et al., 2011), we focus on other three types of compulsions below.
+Inspiration:
+Inspiration: Trigger intensity is affected by not only trigger itself, but also properties of triggers whose variation can result in significantly different anxiety levels.
+As a result, trigger recognition for just-in-time intervention should not only consider object-level detection but also the recognition of certain properties to more accurately predict users’ distress, thus providing suitable intervention.
+4.2. How Triggers Shape Compulsions
+Besides triggers, compulsion is another important indicator to invoke just-in-time interventions.
+Timely detection of compulsive behaviors in daily life is crucial for delivering effective interventions.
+We thus investigated the diverse compulsive behaviors and their relationships to triggers to inspire compulsion detection.
+4.2.1. Distinguishing Compulsive Behaviors
+Compulsive behavior is harder to identify than triggers since compulsions often manifest in ways people with OCD might not even realize, such as avoidance (e.g., avoiding contact with objects deemed dirty by people with contamination OCD) and mental compulsion (e.g., rumination) (T1-T3, T5-T7).
+Therefore, in addition to identifying the typical compulsions (e.g., excessive hand washing) using standardized checklist (e.g., Y-BOCS (Goodman et al., 1989)), professionals would also ask clients about the presence of covert compulsions directly, such as avoidance and mental compulsions (T1-T3, T5-T7).
+In general, our professionals outlined three criteria to help distinguish compulsive behaviors from regular behaviors and assess their impact on daily life:
+  • Functional Impairment (T1): Evaluating how a behavior limits essential activities, such as avoiding social interactions due to intense mental compulsions (e.g., avoiding going outside).
+  • Behavioral Rationale (T5-T7): Determining whether the behavior serves a specific purpose, or if the goal is merely to reduce anxiety (e.g., frequently seeking reassurance from others).
+  • Common Sense (T7): Using common sense to assess whether a behavior is compulsion by considering what others would do to achieve the same goal (e.g., excessive hand washing).
+Inspiration:
+Inspiration: These compulsion identification strategies from professionals could potentially be adapted to support compulsive behavior recognition in real life, for example, with the help of machine learning models that can distinguish common sense from irregular behaviors (Aggarwal et al., 2021).
+4.2.2. Connecting Compulsions to Triggers
+Similar to OCD triggers, our OCD participants experienced diverse compulsive behaviors.
+Unlike prior OCD literature that focused on recognizing compulsive behaviors in itself (e.g., washing hands) (Starcevic et al., 2011; Goodman et al., 1989), we found that our participants constantly associated their compulsive behaviors with triggers.
+Understanding the rationale behind compulsions, particularly how individuals respond to specific triggers, is crucial in interpreting compulsive behaviors, which can potentially help enhance compulsion detection and facilitate intervention delivery at the right timing.
+We thus categorize compulsions into five categories based on their relations to triggers:
 Direct Trigger Interaction.
 Direct trigger interactions are physical behaviors that engage with the trigger to temporarily resolve the distress caused by the trigger.
-For participants with responsibility OCD, repeatedly examining the trigger is a common compulsion, such as checking or jiggling the door locks (P8, P9).
-Repeated checking was also observed in P3 who had contamination OCD.
-She would constantly check the color of the meat to ensure doneness while cooking.
+Repeatedly examining the trigger is a common compulsion for participants with responsibility OCD (e.g., checking or jiggling the door locks, P8, P9), and contamination OCD (e.g., checking the color of meat being cooked to ensure doneness, P3).
 Ordering and organizing is another direct interaction with triggers to precisely control the position, orientation, or order of the triggers to avoid accidents (e.g., keep appliances unplugged when not using, P7, P8), reduce harmful thoughts (e.g., pointing the knife to herself to protect others, P2), and keep life organized and “just right” (e.g., having to park his car in the same spot, P10).
@@
 However, while temporally relieving mental distress, such compulsion could bring long-term mental health issues, “[Hair and eyelash] takes the longest to grow back and it really impacts my confidence and stuff, because it’s noticeable.” (P4)
+Trigger Avoidance.
+Trigger avoidance is another type of compulsions that minimize the interaction with triggers (P1, P3, P6, P8), We found that the extent of avoidance was a spectrum, ranging from partial avoidance, such as rolling their hands into sleeves (P3, P8), or using as few fingers as possible (P6) when touching door knobs, to complete avoidance, such as avoiding eye contact with the trigger (P2).
+Trigger avoidance can lead to severe functional impairment, threatening OCD people’s quality of life.
+P8 expressed his struggle to engage in essential daily activities due to his contamination avoidance, “If there are certain activities that I think are dangerous that’s not safe or something bad could happen.
+It’s gonna get in my head… The biggest challenges are just refusing to engage in a lot of activities to avoid… And I’m not gonna be able to fix the issue the way I want.
+And I’m gonna be stuck somewhere where I can’t do anything about it.
+”
 Consequence Mitigation.
-Consequence mitigation is a compulsive behavior that addresses the speculated aftermath after interacting with a physical trigger when it is inevitable.
+Consequence mitigation is a compulsive behavior that addresses the speculated aftermath after interacting with an object trigger when it is inevitable.
 For participants with contamination OCD, changing/cleaning clothes when coming back home and washing body parts after contacting something ‘dirty’ are common compulsions.
@@
 P3 would feel relieved when someone held doors for her to avoid contamination.
+Mental Compulsions.
+In addition to physical compulsive behaviors, seven participants reported experiencing mental compulsions (P2-5, P7, P9, P10) where they attempted to neutralize the influence of triggers internally.
+For participants with just-right OCD, a common mental compulsion was internal counting, for instance, internally counting delayed time when their schedule was disrupted (P10).
+Other mental compulsions included self-questioning rumination, such as convincing oneself of being a good person despite having intrusive thoughts (P2, P5), catastrophizing about consequences of disrupted routines (P10), and self-reassurance to suppress negative thoughts (P7).
 Effect of Environment Familiarity on Compulsions.
@@
 P4 indicated that she did not want others to see her compulsions, while P6 and P10 ‘gave up’ on compulsions in public knowing that they have no control over the environment.
-In summary, connecting triggers and context information to compulsions can potentially facilitate compulsion recognition by tracking not only user behaviors but also the relationship between users and the triggers or the environment.
-3.2.3. OCD Symptom Management Strategies
-To combat OCD, participants adopted various strategies from OCD therapy (P1-5, P7, P9, P10), support groups (P5), online education resources (P8), and personal experiences (P6, P10), to manage their OCD symptoms in different situations.
-We summarize participants’ coping strategies to inspire the feedback design of in-situ intervention for self-management technology.
-Exposure Practice in Daily Life.
+Inspiration:
+Inspiration: Connecting triggers and context information to compulsions can potentially facilitate compulsion recognition by tracking not only user behaviors but also how specific triggers shape the compulsive reaction of the user.
+As such, a user’s compulsion patterns can be accurately characterized, allowing for more contextualized and personalized intervention design.
+4.3. Managing OCD at Different Stages
+To manage their OCD symptoms, our OCD participants adopted a range of strategies learned from OCD therapy (P1–5, P7, P9, P10), support groups (P5), online educational resources (P8), and personal experiences (P6, P10).
+These strategies were applied at different stages of an OCD episode—from the onset of obsessive thoughts to the execution of compulsions.
+By integrating perspectives from both people with OCD and professionals, we observed substantial overlap in strategy use, but also noted some inconsistencies in how these strategies were applied or recommended.
+We summarize these coping strategies to inform the feedback design of just-in-time interventions for OCD self-management technologies.
+4.3.1. Accepting & Challenging Obsessive Thoughts
+OCD participants employed different mental strategies to address their obsessions, adapted from ACT and cognitive therapy (see Section 2.1).
+Three participants with OCD and nearly all professionals (except T4) used or recommended ACT-based strategies to help shift focus from obsessions to value-driven actions.
+These strategies included: (1) Thought Diffusion (P3, P5, T1, T5)—recognizing and naming the intrusive thought to create psychological distance from the thought; (2) Embracing Uncertainty (P2, P3, T5-T7)—using self-talk like “maybe, maybe not” to accept uncertain outcomes; (3) Refocusing on Values (P5, T1-T3, T5)—identifying values (i.e., important aspects of life) and committing to behaviors aligned with those values.
+While some OCD participants benefited from accepting obsessive thoughts, others managed obsessions by actively challenging and re-evaluating them to reinforce positive thinking (P1, P7), a technique also endorsed by therapists (T3, T5).
+For example, P7 addressed her intrusive fears of unconsciously hitting a pedestrian while driving by evaluating the evidence: “‘okay, what’s the likelihood that this (hitting pedestrians) actually happened?
+And you know, wouldn’t someone else have noticed if this had happened?
+Wouldn’t I have heard, or seen, or touched or smelt things if this had happened?’ So it address like the thought pattern.”
+Inspiration:
+Inspiration: Leveraging the self-reflective nature of these mental strategies, we can design conversation-based persuasive intervention (Orji and Moffatt, 2018) to guide people with OCD to adopt beneficial behaviors and thoughts when obsessions occur, such as intelligent agents that help users navigate their thought processes and address cognitive pitfalls.
+4.3.2. Compulsion Prevention via Real Life Exposure Practices
+When obsessive thoughts were too intense to address or dismiss, participants with OCD resorted to strategies learned from ERP (P1-3, P5, P7, P10) to effectively train their ability to tolerate the anxiety without performing compulsions.
+Planned vs.
+Natural Exposure Practice.
 To improve symptom management ability in a controlled environment, participants would do planned exposure practice as instructed by their therapists (P1-3, P5, P7, P10).
-Specifically, they exposed themselves to a self-created triggering environment at a scheduled time, and resisted compulsions according to the principles of ERP (details in Section 2.1.2), while monitoring their anxiety level.
-P3 articulated her exposure practice targeting the fear of touching batteries, “It (exposure practice) was holding a battery in my hand, rubbing the battery on my arms, putting it in my pocket, rubbing it on my clothes, my phone, on my bag… So I, essentially contaminating all these different things [with batteries] and then, not being able to wash my hands or do anything to decontaminate the battery acid… and having to sit with worrying that everything was contaminated.” Although the flexibility of adjusting trigger intensity allowed participants to gradually build up their tolerance to triggers, most triggers were unplanned in real life, and some triggers that induce intrusive thoughts could not be anticipated (P2).
+In a typical planned exposure practice, people with OCD expose themselves to a self-created triggering environment starting from triggers with moderate intensity on the trigger hierarchy at a scheduled time, and resist compulsions.
+P3 articulated her exposure practice targeting the fear of touching batteries, “It (exposure practice) was holding a battery in my hand, rubbing the battery on my arms, putting it in my pocket, rubbing it on my clothes, my phone, on my bag… So I, essentially contaminating all these different things [with batteries] and then, not being able to wash my hands or do anything to decontaminate the battery acid… and having to sit with worrying that everything was contaminated.” While tolerating the mental distress, people with OCD were instructed to monitor their anxiety level change (e.g., SUDS level) (T2-T6) during the exposure practice.
+In addition to self rating, when possible, professionals detected clients’ anxiety level also by observing their physical behaviors (T3, T7), such as fidgeting, head shaking, tight chest, staring off the trigger, or sometimes an outward expression of distress.
+Typically an exposure practice concludes when their anxiety level decreases (T2, T4, T7).
+Although the flexibility of adjusting trigger intensity allowed participants to gradually build up their tolerance to triggers, most triggers were unplanned in real life, and some triggers that induce intrusive thoughts could not be anticipated (P2, T1, T7).
 Therefore, when confronted with unplanned triggers, participants would seize the opportunity to do a natural exposure practice.
 Due to the spontaneity of unplanned triggers and unpredictable trigger intensity, natural exposure practices were perceived to be more difficult (P1-P3, P7, P10).
-Despite the effectiveness of both planned and natural exposure practices, participants found them challenging especially at the beginning (P2) and tended to avoid exposures when alone (P1).
+Despite the effectiveness of both planned and natural exposure practices, OCD participants found them challenging especially at the beginning (P2) and tended to avoid exposures when alone (P1).
 They had to force themselves to complete the practices, since compulsions were too difficult to resist (P1-P3, P10).
-This reality surfaces a need for in-situ technological support that can assist people with OCD to address their obsessions and compulsions when encountering intense triggers.
-Accepting & Challenging Obsessive Thoughts.
-Participants adopted different mental strategies to address their obsessions, adopted from acceptance and commitment therapy (ACT) (P2, P3, P5, P7-P9) and cognitive therapy (P1, P7, P8) (see Section 2.1.2).
-We summarize strategies used by participants below: (1) Accepting Uncertainty: participants used this strategy to accept the inherent uncertainty in life when their OCD was driven by a need for certainty.
-For example, P3 described her approach of coping with fear of batteries, “Maybe I will die because I am holding this battery.
-Maybe I won’t die.
-I don’t know.
-So a lot of kind of that.
-Maybe this will happen, maybe not.
-And sitting with that uncertainty.” (2) Refocusing on Values: participants identified their values (i.e., important aspects of life) and used them, rather than obsessions, to guide their choice of everyday behaviors.
-(3) Reinforcing Positive Thinking: participants re-evaluated their negative thoughts in a more accurate way to reduce obsessions.
-For example, P7 reframed her intrusive thoughts about unconsciously hitting pedestrians while driving by questioning the presence of evidence, such as any unusual sound.
-Leveraging the self-reflective nature of these mental strategies, we can design conversation-based persuasive intervention (Orji and Moffatt, 2018) to guide people with OCD to adopt beneficial behaviors and thoughts when obsessions occur, such as intelligent agents that help users navigate their thought processes and address cognitive pitfalls.
-Reducing the Urge for Compulsions.
-To suppress the urge of doing compulsions, both mental and physical strategies were used.
-Participants would re-evaluate the necessity of compulsions mentally (P2-4, P6).
-By reflecting on past OCD episodes, participants reminded themselves that nothing bad happened when compulsions were not done (P2) and that compulsions were never working (P4).
+As P10 described, during an exposure practice where the trigger was overwhelming, “it was mentally painful to try and stop (the compulsion).”
+Inspiration:
+Inspiration: These findings highlighted the importance of providing timely and context-aware guidance to help users engage with the exposure rather than avoid it, especially at the beginning, when the user is still building confidence in managing their symptoms.
+Strategies for Resisting Compulsions.
+To better support compulsion prevention, OCD participants reported different mental and physical strategies to suppress the urge of doing compulsions.
+One common strategy is to re-evaluate the necessity of compulsions mentally (P2-4, P6).
+By reflecting on past OCD episodes, participants reminded themselves that nothing bad happened when compulsions were not done (P2) and that compulsions were never working (P4, T1, T7).
 Some weighed the harm compulsions caused against the relief they brought (P3, P4).
@@
 In addition to mental persuasion, participants also employed physical strategies to make compulsions more difficult to act on, e.g., P4 creatively put Vaseline on her eyelash and wore glasses to prevent herself touching her eyelash.
-When all these strategies failed to work, participants (P1, P3-P5, P8, P10) would use self-distraction to temporarily step away from their OCD by switching their attention to casual activities such as watching TV (P3), listening to music (P8), calling a friend (P1), and going out for a walk (P5).
-In contrast to strategies aimed at suppressing the urge for compulsions in the moment, participants with responsibility OCD used reassurance logging to gradually reduce the length and frequency of compulsions (P8, P9).
-They commonly took pictures or videos of their checking behavior to simplify future checking behaviors.
-P8 articulated how reassurance logging alleviated his compulsions, “If I know I’m gonna be gone for a while.
+Completely resisting compulsions overnight can be overwhelming.
+Instead, gradually tapering them—by allowing milder forms of compulsive behavior—can support long-term compulsion reduction.
+Both OCD participants and professionals in our study employed such strategies.
+Participants with responsibility OCD used reassurance logging to gradually reduce the intensity and frequency of their compulsions over time (P8, P9).
+For example, they recorded videos or took photos of their checking behaviors to ease anxiety and avoid the urge to re-check later.
+As P8 explained: “If I know I’m gonna be gone for a while.
 I will take a video where I’ll video my entire apartment, and then I’ll also turn the camera around and say, ‘I have checked the faucet this many times; all plugs are unplugged…’ so it’s almost like telling myself through a video in the moment that I’m good to go.
-And then if I ever get anxious, I can look back at the video.” Although some participants acknowledged that reassurance logging contradicts the principles of ERP (facing their fear), they believed it could complement ERP by gradually building confidence in one’s memory, which could ultimately reduce the need for compulsive checking (P9).
-In summary, participants effectively reduced the urge for compulsions by drawing on past behaviors and memories of previous OCD episodes.
-Inspired by these strategies, symptom and behavior logging could be leveraged to design in-situ intervention methods that increase the cognitive burden of engaging in compulsions.
-Relaxation.
-Unlike above strategies, relaxation techniques were used by participants to reduce the anxiety caused by intense triggers without seeking to break the obsession and compulsion cycle (P1, P3, P6, P7, P10).
-Such strategies included breathing techniques and muscle relaxation.
-Self-Compassion.
-Besides in-situ strategies to combat obsessions, compulsions, and reduce distress, staying motivated is important in symptom self-management.
-When experiencing setback in symptom management that is usually caused by increased trigger intensity, participants would cheer themselves up (P2, P3) and acknowledge their achievements periodically (P5).
-P5 commonly noted down her reflection of success in symptom management—strategies and thought patterns that effectively addressed her obsessions or compulsions.
-External Mental Health Support.
-In addition to self-management strategies, most participants received different forms of support from their family members and friends (P1-8, P10) to help manage their OCD symptoms.
-Some participants had family or friends who understood their mental health conditions and followed therapeutic advice (P2, P4, P7, P10).
-They helped participants monitor obsessions and compulsions to improve self-awareness (P4) and encouraged them to challenge their OCD independently (P1-3, P7).
-In contrast, some participants had family or friends who lacked an accurate understanding of OCD and did not want to see them “suffer.” As a result, they provided accommodation to ‘help’ them avoid triggers, which is usually against therapists’ suggestions.
-For example, P6’s mom washed her hands every time after taking elevators to accommodate P6’s contamination OCD.
-Additionally, participants benefited from support groups (P3, P5, P6) and social media (P1, P8), where people with OCD share their experiences and symptom management strategies.
-To our surprises, none of our participants reported using any technologies dedicated to mental health support.
-3.2.4. Challenges in OCD Symptom Management
-Despite currently available resources and strategies, participants still experience challenges in OCD management.
-OCD-Induced Social Barriers.
-Participants discussed the negative impact of OCD on their social life.
+And then if I ever get anxious, I can look back at the video.” Similarly, professionals allowed clients to engage in reduced compulsions when full exposure practice felt too distressing.
+These included compulsive behaviors with shorter duration (e.g., washing hands for a shorter time, T3, T7) and lower frequency (e.g., checking less often, T2).
+Inspiration:
+Inspiration: In summary, participants effectively reduced the urge for compulsions by drawing on past behaviors and memories of previous OCD episodes.
+Inspired by these strategies, symptom and behavior logging could be leveraged to design just-in-time intervention methods that increase the cognitive and physical burden of engaging in compulsions.
+Tapering strategy can also be translated into just-in-time interventions by prompting users with milder alternatives to their compulsions, helping them gradually build tolerance to anxiety.
+What is Missing in Exposure Practice?
+While all participants were confident about the effectiveness of exposure practices, many acknowledged their limitations in addressing mental compulsions (P2, P5, T1, T2).
+These covert compulsions, such as self-reassurance or repeated mantras, often persisted even when physical compulsions were reduced.
+T1 highlighted this challenge, explaining that some clients might compensate for their behavioral change with mental compulsions, “ They (clients) are doing mental compulsions to sort of cancel out the fact that they didn’t do the physical compulsion, like, okay, I didn’t [check] that doorknob.
+Okay, I’m making my therapist happy, but what I’m doing in my mind is, I’m going over my mantra 15 times, because that makes me safe.” To compensate this limitation, some professionals identified clients’ engagement of mental compulsion alongside their physical compulsive behaviors.
+For example, T2 described asking clients to rate their mental effort during exposure practices to capture these hidden patterns.
+Inspiration:
+Inspiration: This highlights the importance of detecting and monitoring mental compulsions in the design of just-in-time intervention technologies to reveal the full picture of a user’s compulsions.
+4.3.3. Distraction & Relaxation as Safety Net
+When their OCD became so overwhelming that all above strategies failed, OCD participants (P1, P3-P5, P8, P10) turned to self-distraction as a temporary escape—shifting their attention to casual activities such as watching TV (P3), listening to music (P8), calling a friend (P1), or going out for a walk (P5).
+Similarly, relaxation techniques were used to ease anxiety when breaking the obsession–compulsion cycle felt impossible (P1, P3, P6, P7, P10).
+Such strategies included breathing techniques, muscle relaxation, and grounding exercises.
+While professionals supported using these methods to manage intense anxiety (T1, T2), they cautioned that such strategies must be used carefully.
+They noted the importance of helping clients learn to tolerate anxiety rather than eliminate it (T1, T2, T5, T6).
+Without careful guidance, overreliance on distraction or relaxation can itself become a form of compulsion (T1, T2, T5).
+As a result, T7 chose not to teach relaxation techniques at all because they might backfire if they don’t work when clients need them most, “At one point, [the client] may be really distressed.
+And they’ve been practicing these breathing techniques that are not going to work for you, then it’s even more distressing, because the only thing you know how to do [with your OCD] is not gonna work.”
+Inspiration:
+Inspiration: Relaxation strategies can be beneficial to people with OCD only when used correctly.
+It is important for just-in-time technology to involve proper guidance to support and prevent relaxation at the right timing.
+4.3.4. External Mental Health Support
+Psychotherapy is a key resource for learning OCD coping strategies (O’Neill and Feusner, 2015).
+While most participants reported positive experiences, some felt lost and discouraged in therapy not specialized for OCD (e.g., generic talking therapy for P6).
+Nine out of 10 participants also experienced challenges accessing OCD therapy due to long wait time (P1, P5–P7, P9, P10), high costs (P1, P3–P5, P7–P9), and stigma around OCD (P6, P7), underscoring the need for supplementary forms of support.
+Beyond formal therapy treatment, most participants (P1–8, P10) received support from family and friends to help manage their symptoms.
+Some had close contacts who understood their condition and followed therapeutic guidance—helping monitor symptoms (P4) and encouraging them to challenge their OCD independently (P1–3, P7).
+Others, however, had family members or friends who misunderstood OCD and unintentionally reinforced symptoms by providing accommodations contrary to therapists’ suggestions.
+For example, P6’s mother washed her hands every time after using elevators to accommodate P6’s contamination-related fears.
+As several professionals emphasized (T1–5), psychoeducation on how therapy works and the expected outcome of OCD therapy for family members is critical, so they can respond to their loved ones in ways that support, not hinder, treatment progress (T2–4).
+Inspiration:
+Inspiration: Psychoeducation could be integrated into OCD self-management technologies to improve awareness and collaboration between people with OCD and their support system.
+4.4. Opportunities and Gaps in OCD Self-Management Technologies
+In this section, we explore participants’ experiences and attitudes toward existing technologies, and identified key challenges in OCD self-management despite the availability of technological support.
+Based on this, we discuss opportunities to better support OCD self-management.
+4.4.1. Technology Recommendation for OCD Self-Management
+Surprisingly, none of our OCD participants reported using technology specifically designed for mental health support.
+In contrast, OCD professionals recommended a range of tools from symptom tracking to compulsion prevention to help clients manage their OCD outside of therpy (T2-5, T7).
+Dedicated OCD symptom tracking apps, such as OCDfeat (OCDfeat, 2024), were suggested to easily record compulsions and exposure practices (T2).
+They also recommended meditation apps, such as Calm (Calm, 2024), for clients to practice mindfulness (T5) and listen to the bilateral stimulation playlist to calm the nervous system (T4) (Kohl et al., 2014).
+These tools are intended for consistent out-of-session use to monitor symptoms and reinforce skills learned from the therapy.
+Professionals also recommended apps for in-situ use in response to OCD episodes upon triggered.
+To reduce reassurance seeking behavior using specific apps (e.g., web browser), T3 suggested app redirection tool (e.g., ScreenZen (ScreenZen, 2024)) to limit access to those apps on clients’ phones.
+Interestingly, he also recommended text replacement feature to clients to prevent them from searching specific keywords for obsession-related reassurance, “ I’ll have people create shortcuts where… if they’re going to Google ‘heart attack,’ it actually will try to replace ‘heart [attack]’ with like ‘puppies’ or ‘cute cat videos,’ or something like that.”
+Although most participants prefer learning and recommending new technology to their clients, some were reluctant to introduce new tools in the fear of developing reliance on them (T1, T5).
+T1 expressed her concerns, “You want to encourage people to instead learn how to cope with things by themselves… right?
+That’s what we ultimately want.
+So maybe you could have [the technology] where it was only like the first month or the first 3 months, or whatever, and then you have a gradual step down, that might work… because [relying on technology] is counter-therapeutic.
+We want people to be able to help themselves.”
+When talking about self-administered ERP via mobile apps (e.g., NOCD (NOCD, 2024)), professionals recognized its merit in supporting users with low severity (T1-4).
+However, they emphasized the importance of psychoeducation in OCD and acquiring adequate symptom management ability via in-person sessions before using such self-help tools (T1, T3, T4, T6, T7).
+Inspiration:
+Inspiration: To prevent overreliance on technology becoming a new compulsion, just-in-time intervention should adopt a wind-down strategy to gradually help users be independent on symptom management.
+4.4.2. Inconvenience & Privacy Concerns in Symptom Tracking
+Daily symptom tracking helps people with OCD build self-awareness around their behaviors and thoughts (P1, P4, P5, P8).
+It also allows them to observe progress and recognize achievements (e.g., spending less time doing compulsions) over time, which can foster a sense of accomplishment, increase confidence, and strengthen motivation to continue practicing coping strategies (T2–T4, T6).
+Participants typically logged daily triggers (P4, P8), their responses to those triggers (P1, P4), and distress level during exposure practices (P1, P5) either physically or digitally (e.g., Google Doc).
+However, maintaining a consistent tracking habit was challenging (P1, P2, P7, P8, T3, T6).
+OCD participants found traditional journaling methods superficial, as it failed to provide meaningful insights into participants’ symptom patterns or suggest optimal management strategies (P1, P8).
+Professionals, on the other hand, attributed low engagement to forgetfulness due to other daily responsibilities (T3, T6).
+To address this, professionals sometimes had clients fill out symptom tracking forms and set up phone reminders to ensure compliance (T3).
+Privacy concerns also hindered the use of symptom tracking tools (P7, T2).
+Participants were especially cautious about recording sensitive thoughts or exposure practices without clear information on how their data would be stored and protected.
+P7, for example, expressed worry about the emotional risk of recording vulnerable moments: “I would be most concerned about privacy issues, because it’s very vulnerable for people… to do all of that exposure [practices] in a place where their data is [at stake].
+So if sessions are recorded… or their speech or their notes are recorded, making sure all of that is well protected… ” Similarly, T2 noted that clients whose OCD symptoms involved taboo thoughts often avoided logging due to fears about data privacy and access, as she explained, “ They (clients) do worry about putting that (taboo thoughts), even writing it down on paper… so they need a lot of reassurance about where that data is being stored and who has access to it.”
+Inspiration:
+Inspiration: Symptom tracking is a crucial first step for enabling effective just-in-time interventions.
+These findings highlight the need for a simple, transparent tool that can automate tracking while respecting and protecting users’ privacy.
+4.4.3. OCD-Induced Social Barriers
+In addition to challenges with self-management, participants with OCD also described how OCD created difficulties in their social lives.
 Due to the heterogeneity of participants’ OCD symptoms, understanding the nuanced experience of OCD could be difficult, even with OCD professionals (P6, P10).
@@
 Moreover, exposure practices in real life can also bring social barriers since practicing in public spaces or when other people are around could be socially awkward (P1, P7).
-Inconvenient Symptom Logging & Tracking.
-Daily symptom tracking and logging is crucial in helping participants improve self-awareness of their behaviors and thoughts (P1, P4, P5, P8).
-Participants typically recorded daily triggers (P4, P8), their responses to those triggers (P1, P4), and distress level during exposure practices (P1, P5) either physically or digitally (e.g., Google Doc).
-However, participants complained about the limitation of traditional journaling methods since they are unable to provide insights into participants’ symptom patterns.
-When asked about desired technology to close the gap, they proposed an intelligent tracking technology that could recognize their symptom patterns and recommend optimal management strategies based on logging history.
-However, P7 raised concerns over the privacy and emotional risks of such technology, “I would be most concerned about privacy issues, because it’s very vulnerable for people… to do all of that exposure [practices] in a place where their data is [at stake].
-So if sessions are recorded… or their speech or their notes are recorded, making sure all of that is well protected… ”
-3.3. Discussion
-Combating OCD is not an easy task, even with support from various sources.
-While most participants benefited from psychotherapy and developed coping strategies based on their lived experience, tracking and managing OCD outside of therapy sessions can still be challenging.
-Many of these strategies were adapted from psychotherapy, however, without mental health professionals’ feedback, it is unknown whether they were effectively implemented in practices.
-To complement OCD participants’ experiences and guide the design of future intervention technologies, it is crucial to consider OCD professionals’ recommendations for effective symptom management, and further identify gaps between these recommendations and the practices of those living with OCD.
-4. Study 2: Interviewing OCD Therapists
-To gain insights on effective OCD symptom management from OCD experts, we conducted semi-structured interviews with seven mental health professionals who specialized in OCD.
-The study was approved by the Institutional Review Board (IRB) at our university.
-4.1. Method
-4.1.1. Participants.
-We recruited seven OCD professionals (5 female, 2 male).
-Their age ranged from 29 to 60 ($Mean$ = 43.9, $SD$ = 10.4).
-Four participants (T1, T2, T4, T6) had more than 10 years of experiences working as mental health providers.
-Their specialization ranged from OCD, anxiety disorders, depression, eating disorders, PTSD, to substance use disorder.
-Participants had experiences in treating OCD across various severity levels (from mild to severe) and in different practice modalities (both hybrid and virtual).
-Note that the gender imbalance in the participant demographic is consistent with the reality that mental health therapy providers are predominantly women in the US (Clay, 2017; of Labor Statistics, 2023).
-Table 2 shows participants’ detailed information.
-We emailed recruitment information to local OCD therapists listed on International OCD Foundation, a nonprofit organization providing resources and education for people with OCD (IOCDF, 2024).
-Participants were eligible if they were over 18 and had experience conducting OCD related therapy.
-We limited the recruitment to individuals who spoke English and located in the United States.
-Upon completion, each participant received \$40/hour as compensation.
-ID Age/ Gender Certificate Years of Experience Specialization OCD Clients Severity Level Practice Modality
-T1 46/F Licensed Clinical Social Worker 25 years OCD, anxiety disorders, eating disorders, PTSD, depression, substance use disorder Mild to severe Hybrid
-T2 47/F Clinical Psychologist 14 years OCD, axiety disorder Mostly moderate to severe Virtual
-T3 36/M License Professional Counselor 9 years OCD, anxiety disorder, phobias, health anxiety Mild to severe, mostly moderate Virtual
-T4 60/F Licensed Clinical Social Worker 13 years OCD, anxiety, depression Mostly moderate Hybrid
-T5 51/F Professional Counselor Training License (LPC-IT) 2 years and 4 months OCD, anxiety disorder, depression Mild to severe, mostly moderate Hybrid
-T6 38/M License Professional Counselor 17 years OCD, anxiety disorders, eating disorders, PTSD Mild to severe Hybrid
-T7 29/F License Professional Counselor 2 years OCD, anxiety disorder Moderate to severe Virtual
-Table 2. Demographic information and mental health experience of the 7 OCD professionals in Study 2.
-4.1.2. Procedure.
-All interviews were conducted via Zoom.
-Each interview consists of three phases that overall lasted about 90 minutes.
-The first phase focused on participants’ demographic information, certificate, and experience as a mental health care provider.
-In the second phase, we asked about participants’ strategies and technologies used (if applicable) in different treatment stages.
-If a participant conducted ERP, we further asked for details about exposure creation and progress monitoring during exposure practices.
-We then asked about out-of-session homework and external support clients might need.
-In the last phase, we prompted for the challenges of current therapy practices as well as their opinions on current technological support for OCD.
-4.1.3. Data Analysis.
-Similar to Study 1, we conducted thematic analysis on interview transcripts.
-Two researchers coded two representative transcripts (samples) independently and generated a initial codebook.
-One researcher coded the remaining five transcripts using the codebook, updating it with new codes after discussion and agreement with the other researcher.
-Following a hybrid approach (Swain, 2018; Braun and Clarke, 2006), we first generated themes deductively based on therapy treatment process and challenges identified in prior work (e.g., psychoeducation) (Blakey et al., 2017; Wu and Lewin, 2017; Yule and Whittal, 2017).
-We then generate sub-themes and themes (e.g.
-using relaxation techniques suitably) inductively (Boyatzis, 1998).
-The final codebook contained over 300 codes, which we grouped into 16 themes and sub-themes (Table A.3).
-4.2. Findings
-4.2.1. OCD Diagnosis and Assessment
-In the initial assessment of a client’s OCD symptoms, all participants used standardized self symptom rating tools, such as Y-BOCS (Goodman et al., 1989), to gain a basic understanding of their obsessions and compulsions.
-Since compulsions can often manifest in ways clients might not even realize, such as trigger avoidance and mental compulsion (e.g., rumination), our therapists usually conducted interviews to identify the presence of such mental compulsions (T1-T3, T5, T7).
-Due to clients’ lack of awareness of avoidance or mental compulsions, identifying them in early stage is difficult but important for treatment.
-Participants developed criteria to identify compulsive behaviors and assess their impact on daily life:
-  • Functional Impairment (T1): Evaluating how a behavior limits essential activities, such as avoiding social interactions due to intense mental compulsions (e.g., avoiding going outside).
-  • Behavioral Rationale (T5-T7): Determining whether the behavior serves a specific purpose, or if the goal is merely to reduce anxiety (e.g., frequently seeking reassurance from others).
-  • Common Sense (T7): Using common sense to assess whether a behavior is compulsion by considering what others would do to achieve the same goal (e.g., excessive hand washing).
-In addition to in-session assessment, homework was assigned for clients to self-monitor their OCD symptoms now that they have better awareness of their OCD (T1-T3, T5-T7).
-Overall, these strategies could potentially be integrated into self-management technologies to help improve OCD people’s symptoms self-awareness.
-4.2.2. Psychoeducation to Gain Insights
-A person’s insight in OCD refers to the degree to which one recognizes the irrational and excessive nature of their OCD symptoms (Wu and Lewin, 2017), which is considered to be a key factors for successful treatment (T1, T4, T6, T7).
-Clients usually gain insights through psychoeducation along with the initial assessment, where they were taught about the definition of OCD, how OCD therapy works, the expected outcome of the therapy (T1-T7).
-Psychoeducation happens not only didactically but also throughout the treatment process (T1-T4, T6, T7).
-Clients gradually gain insights by discussing their symptoms during therapy sessions and symptom self-tracking between sessions.
-Moreover, to assist with in-session treatment, most participants believed that psychoeducation for clients’ family is equally necessary (T1-5) because clients’ family need to learn how to respond to their loved ones’ behaviors in a therapeutic and helpful way instead of accommodating their OCD (e.g., providing reassurance) (T2-4).
-As such, psychoeducation should also be considered to incorporate into self-management technology for both OCD clients and their families to enhance insights and awareness.
-4.2.3. Active Treatment via ERP
-All participants conducted ERP for their clients.
-In this section, we reported participants’ strategies on identifying trigger intensity, creating exposures, and guiding clients through exposure practices while monitoring their symptoms and progress.
-Building Exposure Hierarchy Using Subjective Experience.
-Participants usually built an exposure hierarchy with clients—a list of OCD triggers ranked based on the intensity of distress they induce—to prepare for the subsequent ERP therapy treatment.
-Along with initial assessment, participants identified triggers via interviews, self symptom rating tools, and the self symptom monitoring homework (T1-T5, T7) and rated the intensity of each trigger.
-Most participants adopted the Subjective Units of Distress Scale (SUDS) (Miegel et al., 2022), a 100-point Likert scale (T1), for self distress assessment.
-Clients were asked to assign each trigger a SUDS score to finalize the exposure hierarchy.
-Uniquely, T3 asked clients to rate their urge to perform compulsions instead of using SUDS to distract them from focusing on the distress associated with triggers, “ If just tracking SUDS of how anxious you feel now, I mean, that’s what people are doing already.
-They’re already hyper vigilant about their anxiety.
-So I need them thinking about something else other than how anxious they feel.”
-Furthermore, similar to our findings in Study 1, some participants examined the properties of triggers that can make a certain trigger less or more intense to build a finer-grained hierarchy (T2, T3, T5), as T2 elucidated, “ There is continuing assessment that’s happening during that hierarchy building.
-Because then I’m asking questions like, what would make [this trigger] easier.
-What would make it harder?
-Maybe they’re just telling me like, ‘well… it depends on if someone’s with me, or if I’m by myself; it depends on how many floors there are in the building…’ whatever variables are important to the exposure.”
-Designing Diverse Exposures.
-Based on the exposure hierarchy, participants designed various types of exposures, including in vivo exposures (T1-T5, T7), imaginal exposures (T2-T7), and material exposures (T2, T3, T5, T6).
-In an in vivo exposure, the client is expected to confront the trigger in real world while supervised by the therapist.
-For exposures that are difficult to create in real world, imaginal exposures were used where participants created a script together with the client and guide them to imagine a scenario.
-Although imaginal exposure is less triggering, it can be effective when used appropriately.
-Two participants (T2, T3) used ChatGPT (OpenAI, 2024) to generate scripts for imaginal exposure.
-For instance, T2 wrote an obituary using ChatGPT for a client who worried about death.
-In addition, material exposures, such as videos, images of triggers, and VR exposure, were typically used to bridge the gap between an imaginal exposure and in vivo exposure.
-Exposure Practices.
-Participants guided clients to participate in exposure practices based on the exposure hierarchy, starting from low intensity triggers.
-Before clients are exposed to triggers, participants first teach them strategies to tolerate anxiety without doing compulsions, aligning with our findings in Study 1.
-During the exposure task, participants would continue to monitor clients’ SUDS level or urge for compulsions (T3) by periodic polling.
-Typically participants ended exposure practices when clients’ SUDS decreases (T2, T4, T7).
-Besides using SUDS, participants detected clients’ distress level by observing their physical behaviors (T3, T7), such as fidgeting, head shaking, tight chest, staring off the trigger, or sometimes an outward expression of distress.
-Whenever these happened, participants brought them up in the discussion after the exposure practice to learn what happened.
-In addition to inhibiting cognitive compulsions, participants also identified clients’ engagement of mental compulsion during the exposure by rating their mental effort (T2).
-During high intensity exposure practices where clients could be too afraid to continue, participants would dynamically adjust the exposure intensity (T1-T7).
-Some participants allowed milder compulsions compared to clients’ usual routines, such as compulsions with shorter duration (e.g., washing hands for shorter time, T3, T7) and lower frequency (e.g., less frequent checking, T2).
-Additionally, participants decomposed the exposure task into simpler steps (T1, T4).
-For example, T1 demonstrated an example where a client avoided opening a new book due to fear of contaminating the book with fingerprints.
-Thus instead of asking them to open the book right away, the client was asked to start with a simpler exposure task—taking the book off the shelf.
-Issues with ERP.
-Overall, all participants were confident about the effectiveness of ERP.
-However, ERP was less effective in managing mental compulsions (T1, T2).
-T1 emphasized the limitation of ERP when clients compensated for their behavioral change with mental compulsions, “ They (clients) are doing mental compulsions to sort of cancel out the fact that they didn’t do the physical compulsion, like, okay, I didn’t [check] that doorknob.
-Okay, I’m making my therapist happy, but what I’m doing in my mind is, I’m going over my mantra 15 times, because that makes me safe.
-”
-Our findings revealed the thorough process of in-session ERP, from building exposure hierarchy, to exposure design and ad-hoc intensity adjustment, providing insights into designing personalized and flexible in-situ technological support for exposure practices.
-4.2.4. Making Long-Term Progress
-To maintain long-term progress, professionals highlighted suitable ways to use certain strategies, assigned homework outside therapy sessions, and recommended self-management technologies.
-Using Relaxation Techniques Suitably.
-Participants agreed that relaxation techniques should be taught to clients with high anxiety level, but advised using them only after exposure practices, as clients are expected to sit with their anxiety rather than resolve it immediately (T1, T2, T5, T6).
-If not used correctly, relaxation techniques can become a new compulsion (T1, T2, T5).
-As a result, T7 chose not to teach relaxation techniques because they might exaggerate clients’ anxiety if they fail to work, “At one point, [the client] may be really distressed.
-And they’ve been practicing these breathing techniques that are not going to work for you, then it’s even more distressing, because the only thing you know how to do [with your OCD] is not gonna work.”
-Promoting Out-of-Session Success via Homework.
-Homework is crucial for successful OCD treatment.
-Participants assigned homework for clients to practice both planned and natural exposures (T1-7) outside of sessions (Section 3.2.3), where clients were asked to record their compulsion duration, anxiety level, and time needed for anxiety to decrease during each practice (T2-T7).
-Along with exposure practices, clients were instructed to implement learned strategies to promote cognitive change, such as negative thought reframing (T5, T7), ACT-related strategies (e.g., uncertainty mindset) (T1-3, T5-7), which overlap with self-management strategies mentioned in Study 1.
-By tracking their compulsive behavior and anxiety level, clients became more aware of their progress and achievements across therapy sessions (e.g., spending less time washing hands).
-Such positive reflections on progress can enhance confidence and perseverance (T2-4, T6).
-However, participants also recognized that self-tracking can be difficult to maintain, due to forgetfulness and other daily responsibilities (T3, T6).
-In addition, exposure practice itself can be overwhelming with multiple triggers present out-of-session, leading to low homework compliance (T3, T7).
-To address these issues, participants would identify specific barriers faced by individual client, such as fear of exposure (T1, T3), and adjust the homework content accordingly.
-Although monitoring clients’ progress seemed impossible out-of-session, participants sometimes had clients fill out exposure tracking forms and set up phone reminders to ensure homework completion (T3).
-Technology Recommendation for Self-Symptom Management.
-To ensure clients implementing therapeutic advice outside of therapy, participants also recommended tools and technologies to help clients manage OCD at home (T2-5, T7).
-Dedicated OCD symptom tracking apps, such as OCDfeat (OCDfeat, 2024), were suggested to easily record compulsions and exposure homework (T2).
-They also recommended meditation apps, such as Calm (Calm, 2024), for clients to practice mindfulness (T5) and listen to the bilateral stimulation playlist to calm the nervous system (T4) (Kohl et al., 2014).
-These tools are intended for consistent out-of-session use to monitor symptoms and reinforce skills learned from the therapy.
-Our participants also recommended apps for in-situ use in response to OCD episodes upon triggered.
-To reduce reassurance seeking behavior using specific apps (e.g., web browser), T3 suggested app redirection tool (e.g., ScreenZen (ScreenZen, 2024)) to limit access to those apps on clients’ phones.
-Interestingly, he also recommended text replacement feature to clients to prevent them from searching specific keywords for obsession-related reassurance, “ I’ll go in and have people create shortcuts where… if they’re going to Google ‘heart attack,’ it actually will try to replace ‘heart [attack]’ with like ‘puppies’ or ‘cute cat videos,’ or something like that.”
-However, participants found low engagement in recommended technologies among clients, especially for tracking purpose (T2, T3).
-While T3 attributed the issue to poor usability of the outdated interface, T2 identified privacy concerns for clients whose OCD symptoms involve taboo thoughts.
-As she explained “ They (clients) do worry about putting that (taboo thoughts), even writing it down on paper… so they need a lot of reassurance about where that data is being stored and who has access to it.”
-Although most participants prefer learning and recommending new technology to their clients, some were reluctant to introduce new tools in the fear of developing reliance on them (T1, T5).
-T1 expressed her concerns, “You want to encourage people to instead learn how to cope with things by themselves… right?
-That’s what we ultimately want.
-So maybe you could have [the technology] where it was only like the first month or the first 3 months, or whatever, and then you have a gradual step down, that might work… because [relying on technology] is counter-therapeutic.
-We want people to be able to help themselves.”
-When talking about self-administered ERP via mobile apps (e.g., NOCD (NOCD, 2024)), participants recognized its merit in supporting users with low severity (T1-4).
-However, participants emphasized the importance of psychoeducation in OCD and acquiring adequate symptom management ability via in-person sessions before using such self-help tools (T1, T3, T4, T6, T7).
-In general, participants highlighted the need for a tracking tool that preserves privacy.
-Inspired by participants’ strategies to facilitate long-term success, future self-management technology should guide users to choose strategies wisely, and incorporate progress reflection to maintain high motivation.
+Inspiration:
+Inspiration: More support is needed across social, clinical, and technological contexts to foster greater awareness of OCD and ensure that people with OCD have access to supportive environments and resources.
 5. Discussion
-As an important part of treatment process, self-management of OCD presents significant challenges.
-The needs for technological support has long been under explored.
-Through our two studies with people with OCD and the OCD professionals, we sought to triangulate the current gaps and opportunities in OCD symptom self-management.
-From both studies, we found that people with OCD employed strategies learned from mainstream evidence-based therapies such as ERP and ACT and developed experience-based strategies to resist the urge of compulsions, aligning with professionals’ advice.
-This confirms the efficacy and generalizability of strategies in OCD therapy.
-However, the two studies also highlighted gaps in OCD self-management outside of therapy despite currently-available resources.
-First, while natural exposure practice is encouraged due to their abundant presence in people’s daily life, these practices were often difficult to complete and tended to be avoided, confirming prior research (Seibell and Hollander, 2014; Eddy et al., 2004).
+As an important aspect of quality of life, self-management of OCD presents significant challenges and the needs for technological support has long been underexplored.
+Our research fill this gap by thoroughly investigating the perspectives of both people with OCD and OCD professionals.
+Our findings centered on the lived experiences of people with OCD and supplemented and contrasted with the professionals’ views on treatment strategies and technology recommendations.
+As a result, we identified the gaps and opportunities in OCD symptom self-management.
+First, while natural exposure practice was encouraged due to their abundant presence in daily life, these practices were often difficult to complete and tended to be avoided (Seibell and Hollander, 2014; Eddy et al., 2004).
 Besides the difficulty of facing the fear, identifying mental compulsions could be challenging and required adequate insights.
 The absence of physical compulsion does not guarantee the absence of mental compulsions despite going through ERP.
-Second, some experience-based strategies adopted by people with OCD (Study 1) were identified to be counter-therapeutic and suggested to be prohibited (e.g.
-reassurance) or used cautiously (e.g., relaxation techniques) by the professionals due to the potential of developing reliance (Study 2).
-Third, both OCD clients and therapists recognized the importance of symptom and activity tracking, however, traditional tracking tools present limitations and raise privacy concerns due to vulnerability and shame associated with OCD, resulting in the low technology adoption in Study 1 and echoing OCD therapists’ observations in Study 2.
-Finally, people with OCD faced persisting challenges in OCD self-management despite receiving therapy, highlighting the need for additional support to facilitate smooth social interactions.
-The insights from both studies emphasized the needs for self-management technologies beyond formal therapy sessions.
-By triangulating the insights from people with OCD and the professionals, we identify diverse OCD triggers and properties, compulsive behaviors in relation to triggers, and various effective strategies to inspire the input (e.g., trigger and compulsion recognition) and output (i.e., feedback design) for in-situ self-management technologies for OCD.
-Furthermore, given the persisting challenges experienced by OCD people, we discuss how technology can address OCD-specific privacy issues and opportunities to support OCD people’s social needs.
+Second, some experience-based strategies adopted by people with OCD (e.g., relaxation techniques) were suggested to be used cautiously by the professionals due to the potential risk of developing reliance.
+Third, both OCD participants and therapists recognized the importance of symptom and activity tracking, however, traditional tracking tools present limitations and raise privacy concerns due to vulnerability and shame associated with OCD, resulting in low technology adoption.
+Finally, people with OCD faced persisting challenges in OCD self-management especially in social scenarios despite receiving therapy, highlighting the need for additional support to facilitate smooth social interactions.
+These gaps emphasized the needs for self-management technologies beyond formal therapy sessions.
+Based on the diverse OCD triggers, compulsive behaviors, and symptom management strategies identified in our research, we discuss the input (e.g., trigger and compulsion recognition) and output (i.e., feedback design) of future just-in-time self-management technologies for OCD.
+Furthermore, given OCD people’s challenges, we discuss how technology can address OCD-specific privacy issues and opportunities to support OCD people’s social needs.
 5.1. Enhancing Symptom Recognition & Tracking with User Input
-As an important step of insight establishing and symptom self-management, our studies underscored a need for supporting personalized, automatic, ad fine-grained tracking of OCD symptoms and exposure practices.
+As an important step of establishing symptom self-awareness for symptom self-management, our study underscored a need for supporting personalized, automatic, and fine-grained tracking of OCD symptoms and exposure practices.
 5.1.1. Supporting Personalized & Automatic OCD Tracking
-From Study 1, we found that OCD triggers are highly individualized, taking diverse forms (e.g., objects, contexts, sounds, thoughts), and that their properties significantly influence trigger intensity.
-While symptom tracking is essential for improving insights, it is difficult to maintain without therapist supervision (Study 2).
+We found that OCD triggers are highly individualized, taking diverse forms (e.g., objects, contexts, sounds, thoughts), and that their properties significantly influence trigger intensity.
+While symptom tracking is essential for improving symptom awareness and motivation, it is difficult to maintain without therapist supervision.
 Therefore, future technology should incorporate OCD users’ input, allowing specification of triggers to support personalized and automatic trigger detection.
-Similarly, the system can prompt for the user’s compulsive behaviors based on different types of compulsions (Section 3.2.2) and criteria proposed by OCD therapists (Section 4.2.1).
+Similarly, the system can prompt for the user’s compulsive behaviors based on different types of compulsions and criteria proposed by OCD therapists (Section 4.2).
 Recent advances in AI have made it possible to adapt pre-trained models to specific domains through few-shot learning (Parnami and Lee, 2022; Zhang et al., 2023).
 By leveraging just a few examples provided by users, these AI models can be personalized to effectively detect and track OCD symptoms within each user’s unique context.
-Given the varied form of triggers and compulsions (e.g., direct trigger interaction, consequence mitigation), multi-modal AI models (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024) can be used together with wearable devices (e.g., egocentric camera and microphone (Liu et al., 2024)) to identify and track triggers and compulsions beyond the scope of visual-based methods.
+Given the varied form of triggers and compulsions (e.g., direct trigger interaction, consequence mitigation), multi-modal AI models (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024) can be used together with wearable devices (e.g., egocentric camera and microphone (Liu et al., 2024)) to identify and track triggers and compulsions beyond the scope of vision-based methods.
 Despite the merit of AI-based symptom tracking automation, false positive prediction results can potentially exacerbate OCD symptoms, since people with OCD often have low confidence in their memories (Radomsky et al., 2014).
@@
 Besides highly personalized triggers, our findings suggest the importance of considering different properties of triggers to enable fine-grained exposure hierarchy building and symptom tracking.
-In complement to personalized trigger specification, future tracking technology should prompt the user to reflect on how trigger intensity is affected by different properties of that trigger (Section 3.2.1), measured with SUDS or urge to perform compulsion (Section 4.2.3).
+In complement to personalized trigger specification, future tracking technology should prompt the user to reflect on how trigger intensity is affected by different properties of that trigger, measured with SUDS or urge to perform compulsion (Section 4.1).
 As such, an exposure hierarchy with fine-grained and structured trigger properties can be built.
 During tracking, the system can log fine-grained details of triggers and situate them in the hierarchy as well as identify physical compulsive behaviors, to better support the tracking of exposure practices.
-Figure 1a shows a usage scenario that can be achieved with proposed system: the user identified knife as a trigger of his harm OCD, they then are prompted to answer how each property can affect the trigger intensity, for example, a knife can be more triggering when the knife is closer to him, so they assign ‘knife in hand’ with the highest intensity level, and ‘knife on countertop’ with a moderate intensity level.
-They are also prompted to describe their compulsion—‘avoiding the knife.’ The system will then detect the physical triggers and compulsions in real-time from the camera feed.
-For mental compulsions which are difficult to recognize, potential technology can incorporate physiological sensing (e.g., EEG, ECG, eye movement) to detect mental stress (Xia et al., 2019; Al-Jumaily et al., 2021) and mental workload (Fan et al., 2020; Wang et al., 2023), inspired by OCD therapists’ strategies in Study 2.
-Prior research has shown great potential of using physiological sensing as a therapeutic approach to detect users’ internal states for various mental health conditions.
-For example, estimating the anxiety level during exposure therapy for anxiety disorder (Gazi et al., 2021; Šalkevicius et al., 2019), and detecting the presence of depression to aid clinical diagnosis(Saeedi et al., 2021; Ay et al., 2019; Thoduparambil et al., 2020).
-These methods can be adopted to support OCD self-management.
-However, context information should be incorporated with physiological data since both our work and prior literature (Llorens-Aguilar et al., 2021) indicated the importance of context in recognizing OCD symptoms.
-Overall, the proposed technology can potentially improve tracking compliance and precision by formulating fine-grained exposure hierarchy through natural language, and automating OCD symptom tracking using state-of-the-art AI systems.
-However, we do recognize that our proposed technology could not capture the same level of detail for tracking mental triggers (e.g., worries).
-5.2. In-Situ Feedback Design for OCD Self-Management
+Figure 1 shows a usage scenario that can be achieved with proposed system: the user identified “trash can” as a trigger of their contamination OCD, they then are prompted to answer how each property can affect the trigger intensity, for example, a trash can can be more triggering when the trash can is overflowing, so they assign “full of trash” with the highest intensity level, and “no trash” with a moderate intensity level.
+They are also prompted to describe their compulsion—“avoiding using the trash can.” The system will then detect the physical triggers and compulsions in real-time from the camera feed.
+Mental compulsions are often difficult to recognize but critical to address.
+To support their detection, future technologies could incorporate physiological sensing, such as EEG and ECG, to assess mental workload (Fan et al., 2020; Wang et al., 2023), inspired by strategies used by professionals (T2).
+Prior research has also shown promising results in detecting mental rumination using self-report measures combined with wearable sensor data, like EMG (Nalborczyk et al., 2017; Arnold and Young, 2025).
+By capturing a user’s full compulsion pattern and integrating contextual information (e.g., triggers, environment), the system can generate more tailored and effective intervention strategies.
+Despite the potential, prompting users to report symptoms itself can be triggering.
+Our suggested input methods may require a certain level of symptom awareness and readiness for symptom management, which should be assessed before prompting users’ symptoms.
+Addressing this challenge is essential for creating safe and effective self-management tools.
+Future research should explore symptom reporting methods that minimize discomfort while ensuring honesty.
+Figure 1. Fine-Grained Symptom Formulation: an example of the user specifying triggers and trigger properties to build fine-grained exposure hierarchy.
+[…]
+The figure mainly contains three part left to right.
+On the left, it shows a conversation interface between a user with OCD and an AI agent.
+The agent asks the user ‘What is your OCD trigger?’ the user answers ‘Trash cans.’ Then the agent says ‘Please tell me how following properties of a knife affect its trigger intensity.
+Let’s start with Size.’ The user’s speech bubble shows ‘…’ In the middle, there are three sliders labeled with different properties.
+All three sliders are of same length with the left end representing ‘Lowest SUDS’, and right end representing ‘Highest SUDS’.
+There are two knobs on the ‘size’ slider saying ”wastebasket at home” and ”big garbage can outside” located at the middle-left and the right end of the slider, respectively.
+There are two knobs on the ‘Trash inside?’ slider saying ”no trash” (located near the left end of the slider) and ”full of trash” (located on the right end).
+There are two knobs on the ‘With a lid?’ slider saying ”no lid” (located near the left end of the slider) and ”lid with a handle” (located near the right end).
+On the right of the figure, it shows an illustration of a silver trash can with a recycling symbol on it.
+5.2. Just-in-Time Feedback Design for OCD Self-Management
 Consistent with prior literature (Lind et al., 2013; Hezel and Simpson, 2019; Ojalehto et al., 2020), exposure practices can be difficult to follow through outside of therapy sessions, due to the existence of fearful triggers in diverse context.
-Triggers can appear anywhere anytime and provoke users’ OCD spontaneously, the burden of dealing with unpredicted triggers and resisting compulsions thus all fall onto the users with OCD, potentially resulting in severe consequences.
-We thus propose the following technologies that provide in-situ intervention during high intensity exposures in people’s everyday life.
-5.2.1. In-Situ Compulsion Prevention
-Inspired by participants’ strategies in Study 1, and OCD therapists’ technology recommendation, future intervention technology should consider using restrictive interactions to make compulsions more difficult to perform.
-In the field of human-computer interaction (HCI), researchers have explored the design of such techniques to prevent smartphone overuse (Orzikulova et al., 2024; Lu et al., 2024; Xu et al., 2022), such as delaying a user’s tap for a certain period before it takes effect (Lu et al., 2024).
-In OCD context, these restrictive interactions can be adapted to help reducing compulsions in digital space, e.g., seeking reassurance online.
-For compulsive behaviors that happen in physical world, potential technology can explore the design of mixed reality (MR) -based interaction to suppress the urge for compulsions.
-Researchers has explored the usage of virtual barriers to ensure user safety in MR (Hoang et al., 2022; Sousa et al., 2019).
+Triggers can appear anywhere anytime and provoke users’ OCD spontaneously.
+As such, the burden of dealing with unpredictable triggers and resisting compulsions all fall onto the users with OCD.
+This can potentially result in severe consequences.
+We thus propose the following technologies that provide just-in-time intervention during high intensity exposures in people’s everyday life.
+5.2.1. Augmented Reality-Based Compulsion Prevention Support
+Inspired by OCD participants’ strategies and OCD therapists’ technology recommendation, future intervention technology should consider using restrictive interactions to make compulsions more difficult to perform.
+Inspired by prior work on smartphone overuse prevention (Orzikulova et al., 2024; Lu et al., 2024; Xu et al., 2022), future intervention technology can adopt similar restrictive interactions to help reducing compulsions in digital space, e.g., seeking reassurance online.
+For compulsive behaviors that happen in physical world, potential technology can explore the design of augmented reality (AR) -based interaction to suppress the urge for compulsions.
+Researchers has explored the usage of visual barriers in AR and VR (virtual reality) systems to ensure user safety (Hoang et al., 2022; Sousa et al., 2019).
 Such design can be adapted for compulsion prevention.
-Since most overt compulsions involve visual examining, disturbing the user’s vision can effectively interrupt compulsions.
-Figure 1.
-(a) Fine-Grained Symptom Formulation: an example of the user specifying triggers and trigger properties to build fine-grained exposure hierarchy; (b) AR Compulsion Intervention: an example of using visual augmentation (blurry effect) to disturb the user’s vision.
-Figure 1b describes a scenario where the user with contamination OCD washes their hands after contacting triggers (e.g., dirty dishes).
-Once the system detects excessive hand washing behaviors (longer than 20s), it will blur out or render distracting patterns on the users’ hands and faucet — objects involved in hand washing compulsions.
-As such, the user’s vision becomes too low to examine their hands, and eventually stopped washing them.
-Adopting the idea of allowing milder compulsion (Section 4.2.3) when exposure becomes too intense, the invasiveness (e.g., blurry effect intensity) of the visual augmentation should be adjustable.
-Such intervention can be promising for direct trigger interaction and consequence mitigation that involves active interaction between the user and surrounding physical objects.
-For trigger avoidance type of compulsions, such as avoiding eye contact with people that induce intrusive thoughts, the system can render visual augmentations that encourages interactions with the trigger.
+Since most overt compulsions involve visual examining (e.g., repeated checking), adequately disturbing the user’s vision by placing visual barriers on objects being examined may effectively interrupt physical compulsions.
+For instance, rendering visual blockage on stovetop when the user repeatedly checking burner knobs.
+For trigger avoidance type of compulsions, such as avoiding eye contact with people that induce intrusive thoughts (P2), the system can render visual augmentations that encourages interactions with the trigger.
 For instance, when avoidance of triggers is detected, the system can render a gray-scale filter on everything the user sees except the triggering person to encourage eye contact, and turn off the filter when eye contact is maintained for a preset length of time as an exposure practice.
@@
 Future work should explore the optimal balance between effectiveness and restrictiveness of such technologies in OCD context.
-5.2.2. In-Situ Cognitive Support
-Besides physical methods, mental support inspired by ACT and cognitive therapy is another important strategy that effectively adjust people’s attitudes on their obsessions.
-Self-managment technologies should thus consider employing contextualized suggestions to facilitate acceptance of thoughts, once mental distress or compulsion is detected.
-For instance, the technology can guide the user to acknowledge the uncertainty to the user’s feared consequences after encountering the trigger.
-Combined with professional expertise, recent advancement in LLMs have enabled such context-aware intervention to support general mental well-being (Kumar et al., 2024; Nie et al., 2024).
-Moreover, value-based intervention has been suggested and implemented in prior work (Khan et al., 2011; Xu et al., 2022) as a ‘wake up call’ to remind the user of committing to their values.
-For example, when the user with contamination OCD is compulsively washing hands, a photo of their family will be displayed on their phone to encourage them to stop, if they identified ’spending time with family’ as a value.
-Alternatively, we can design in-situ cognitive support based on cognitive therapy.
-Such technology can guide the user via conversations to challenge their obsessions, such as reframing obsessive thoughts to more hopeful ones (Sharma et al., 2023), and analyzing the likelihood of feared consequences to reject obsessive thoughts (Yule and Whittal, 2017).
-Although the feedback design was proposed for in-situ self-management, we believe they can be effectively adopted for out-of-session practices to reinforce therapy outcome.
+5.2.2. LLM-based Cognitive Support
+In addition to physical methods, mental strategies inspired by ACT and cognitive therapy are essential for helping individuals reshape their attitudes toward obsessions.
+LLMs can be leveraged to offer personalized suggestions based on these strategies when mental distress or compulsions are detected.
+For instance, the technology can guide the user to embrace uncertainty following a triggering event.
+Moreover, value-based intervention techniques (Khan et al., 2011; Xu et al., 2022) can serve as a ‘wake up call’ to remind the user of committing to their values—for instance, displaying a photo of a loved one on the user’s phone to interrupt compulsive hand-washing, if the user has identified “spending time with family” as a core value.
+Alternatively, we can design intervention based on thought reframing—guiding the user via conversations to challenge their obsessive thoughts and reassess the likelihood of feared outcomes (Sharma et al., 2023; Yule and Whittal, 2017).
+However, generating effective, context-aware responses requires carefully crafted prompts, which can be difficult without domain-specific expertise (Priyadarshana et al., 2024).
+Although prior work has explored LLM-based support for general CBT (Kian et al., 2024; Huang et al., 2024), no existing study has focused on adapting these techniques for OCD-specific contexts.
+Future work should explore not only prompt engineering for pre-trained LLMs but other techniques such as retrieval-augmented generation (Lewis et al., 2020; Guo et al., 2024) to integrate OCD-specific knowledge and user history for more targeted and effective intervention.
 5.3. Addressing Privacy Needs for People with OCD
@@
 However, due to technology limitation, many AI services require sending user data to remote servers for inference (OpenAI, 2024).
-In the context of OCD, researchers should pay attention to sensitive information that can exacerbates OCD users’ mental health issues.
+In the context of OCD, researchers should pay attention to sensitive information that can exacerbates OCD users’ mental health conditions.
 Users should be prompted to identify the type of sensitive data, such as audio or video involving fear, and taboo thoughts (e.g., sexual and violent thoughts).
@@
 5.4. Reducing Frictions in Social Interactions
-Our studies suggest that OCD poses challenge to smooth social interactions due to the need for compulsions.
+Our study suggest that OCD poses challenge to smooth social interactions due to the need for compulsions.
 Prior work also revealed that people with OCD experienced more interpersonal problems primarily due to certain worries and rumination than people without OCD (Solem et al., 2015).
 Due to the low public awareness of OCD, people with OCD usually need to endure the awkwardness caused by their symptoms during social interactions.
-Disclosing one’s OCD status can earn accommodation or understanding from people (P10 in Study 1).
-However, not all people with OCD are willing to disclose their mental health status, particularly when their OCD involves taboo thoughts (e.g., P2’s sexual OCD in Study 1).
+Disclosing one’s OCD status can earn accommodation or understanding from people (P10).
+However, not all people with OCD are willing to disclose their mental health status, particularly when their OCD involves taboo thoughts (e.g., P2’s sexual OCD).
 For those who are still building symptom management skills, additional support is needed to reduce frictions caused by OCD in everyday social interactions.
 With current technology, researchers can explore the use of motivational interviewing via conversational agent (Miller and Rollnick, 2012; Oewel et al., 2024) to encourage the user to keep engaged with social interactions.
-Technology that delivers unobtrusive intervention for emotion regulation can also be used to address the mental distress in such situation, e.g., reducing anxiety level via haptic feedback simulating a low heart rate (Costa et al., 2016, 2019; Kelling et al., 2016).
+Technology that delivers unobtrusive intervention for emotion regulation can also be used to address the mental distress in such situation (Costa et al., 2016, 2019; Kelling et al., 2016).
 That said, technology can provide limited assistance to satisfy OCD people’s social needs without improved public awareness of OCD, and appropriate social support.
@@
 5.5. Technology Adoption Concerns
-The low adoption rate of technology by people with OCD revealed in both studies raises concerns about the practicality of our proposed technology design opportunities.
+The low adoption rate of technology by people with OCD revealed in our study raises concerns about the practicality of our proposed technology design opportunities.
 Prior research in accessibility has pointed out key factors that affect the adoption of assistive technologies, including ease of use (Scherer, 2017; Phillips and Zhao, 1993), trust (Scherer, 2017), self-consciousness (Parette and Scherer, 2004; Scherer, 2017), and social acceptance (Shinohara and Wobbrock, 2011; Scherer, 2017).
-These factors all can potentially affect the adoption rate of suggested technology for OCD in-situ intervention.
+These factors all can potentially affect the adoption rate of suggested technology for OCD just-in-time intervention.
 In addition, mixed-reality is still in its early stages (Cummings et al., 2023), and options for consumer level physiological sensing technology is limited.
@@
 Our research has several limitations.
-First, while participants in Study 1 represented diverse OCD themes, it is unclear whether our findings apply to less common OCD themes, e.g., relationship OCD.
-Future work should include a broader range of OCD themes to explore differences in management strategies and technology needs.
-Second, we recruited participants with mild to moderate OCD at the time of the study (Study 1).
-Although nine had experienced severe OCD before therapy, their self-management strategies and needs may vary with symptom severity.
+While our OCD participants represented diverse OCD themes, it is unclear whether our findings apply to less common OCD themes, e.g., relationship OCD.
+Future research should include a broader range of OCD themes to explore differences in management strategies and needs.
+In addition, we recruited participants with mild to moderate OCD at the time of the study.
+Although nine had previously experienced severe OCD symptoms, their self-management strategies and needs may vary with symptom severity.
 Hence, our findings should be generalized to people with severe OCD with caution.
+Although interviews with therapists enriched our understanding of treatment frameworks and common challenges in symptom management, their perspectives reflect clinical knowledge rather than lived experience.
+While prioritizing the views of people with OCD, the involvement of OCD professionals in our study may have influenced how we organized and interpreted participant responses.
+Future research should involve more people with OCD from diverse background as co-researchers to ensure that lived experience remains central throughout the research process and to counterbalance clinical assumptions.
+We acknowledge the potential compensation issue in our study.
+We provided reasonable compensation to both participant groups but offered OCD professionals a higher hourly rate to facilitate recruitment given their limited availability.
+This imbalance in compensation may raise ethical concerns.
+Future research should strive to provide comparable compensation for similar tasks across participant groups, based on time commitment and participation burden, to ensure fairness in the recruitment process.
+Lastly, our work is a need-finding study (Patnaik and Becker, 1999) aimed at surfacing challenges in OCD self-management and identifying technology opportunities.
+In the future work, we will implement and evaluate the interventions proposed in this study to validate their feasibility and effectiveness.
 7. Conclusion
-Our research contributed the first in-depth investigation to understand OCD people’s specific experiences and their technology needs for symptom self-management.
-Through an interview study with 10 participants with diverse OCD symptoms, we identified their individualized triggers, compulsions, properties that affect trigger intensity, and revealed their self-management strategies.
-With another interview study with seven OCD therapists, we revealed the nuances in treatment strategies in complement to those used by people with OCD.
-Combining the two studies, we identified gaps in OCD self-management despite currently available resources, and discuss design opportunities for personalized symptom tracking, in-situ OCD intervention, and addressing privacy and social needs that are unique to OCD.
+Our research contributed the first in-depth investigation to understand OCD people’s specific experiences and their technology needs for just-in-time symptom management.
+Through an interview study with 10 participants with diverse OCD symptoms and seven OCD professionals, we identified individualized triggers and compulsions, explored properties that influence trigger intensity, and discussed how these symptoms can be recognized and tracked.
+We also examined the self-management strategies adopted by OCD participants, highlighting both alignment with and deviations from professional-recommended practices.
+Finally, we identified gaps in OCD self-management despite currently available resources, and discuss design opportunities for personalized symptom tracking, just-in-time OCD intervention, and addressing privacy and social needs that are unique to OCD.
+Acknowledgements.
+We sincerely thank all participants who generously shared their time, experiences, and insights with us.
+This work would not have been possible without their willingness to contribute.
+We also thank Ting Zhou and Linda Zeng for their valuable input on the study design during the early stages of the project.
 Appendix A Appendix
@@
 Table A.1.
-Detailed personal experience of OCD (obsessions, triggers and compulsions) of reported by participants in Study 1.
-Only representative triggers are listed in cases where participants reported more than three triggers.
-Themes Sub-themes Example codes
-Trigger characterization Spectrum of triggers sound of utensil bumping, before bed time, doubt
-Sensory attributes color, sticky texture, liquid more gross than solid
-Affordances door knob, bar in bus, trash can with a lid
-Spatial attributes proximity to trigger, amount of triggers, position/pose of trigger
-Certainty of contamination origins unknown food source, who touched the food, people they trust
-Other factors affecting trigger intensity mental stress, self-care, physical stress
-Connecting triggers to compulsions Direct trigger interaction jiggling the door lock, moving stuff from heating vents
-Consequence mitigation cleaning, changing outfit, body scanning for symptoms
-Assisted compulsions seeking reassurance, asking whether telling truth
-Effect of environment familiarity on compulsions more intense with closed people, compulsion less prominent in public
-OCD symptom management strategies Exposure practice in daily life contaminate with battery, maintaining eye contact with triggers
-Accepting & challenging obsessive thoughts uncertainty mindset, thought diffusion, focusing on values
-Reducing the urge for compulsions self reminding compulsion is unnecessary, compulsion not working from experience, make compulsion more difficult
-Reassurance logging taking picture, changing to 24hr time
-Relaxation breathing skills, muscle relaxation
-Self-compassion reflections from success, celebrate breakthrough day
-External mental health support not judging shameful behaviors, family accommodating to OCD
-Challenges in OCD symptom management OCD-induced social barriers OCD interrupt social interactions, downplay to get trust
-Inconvenient symptom logging & tracking no self-tracking, digital journal on symptom tracking
-Table A.2. Themes, Sub-themes and Example Codes generated from Study 1
-Themes Sub-themes Example codes
-OCD diagnosis and assessment Identify triggers interview, self-monitor triggering events
-Identify compulsion untypical compulsion, mental compulsion, duration on compulsion
-Psychoeducation to gain insights Psychoeducation as standalone session explain how therapy works, explain what to expect
-Improve behavior awareness by homework self-monitor intrusive thoughts, self-monitor compulsion
-Active treatment via ERP Building exposure hierarchy using subjective experience SUDS, ask extent to stay in anxiety, ask extent to conduct rituals
-Examine trigger properties identify fear together, fluid process, identify wiggle room
-Designing diverse exposures in vivo exposures, imaginal exposures, material exposures
-Exposure Practices distress detection, ask mental compulsion, practice until decreased SUDS
-Issues with ERP only focus on behavioral change, mental compulsion to compensate behavioral
-Making long-term progress Using relaxation techniques suitably only after exposure, only for long exposure, relaxation causes hopelessness
-Promoting out-of-session success via homework people not good at tracking , identify homework barriers, afraid to do homework
-Technology recommendation for self-symptom management OCDefeat, low motivation, mindfulness app, concern about privacy
-Table A.3. Themes, Sub-themes and Example Codes generated from Study 2
+Detailed personal experience of OCD (obsessions, triggers and compulsions) reported by participants.
+For participants who reported more than three triggers, only representative examples are listed.